### v2

# Curvature and local matchings of conference graphs and extensions

Abstract
We confirm a conjecture of Bonini et al. on the precise Lin–Lu–Yau curvature values of conference graphs, i.e., strongly regular graphs with parameters $(4\gamma+1,2\gamma,\gamma-1,\gamma)$, with $\gamma\geq 2$. Our method only depends on the parameter relations and applies to more general classes of amply regular graphs. In particular, we develop a new combinatorial method for showing the existence of local perfect matchings. A key observation is that counting common neighbors leads to useful quadratic polynomials. Our result also leads to an interesting number theoretic consequence on quadratic residues.
1 Introduction
The study of discrete notions of Ricci curvature has attracted much attention and found important applications in graph theory, see, e.g., [2, 3, 10, 21, 23, 24, 28, 29]. Ricci curvature is a key concept in Riemannian geometry and geometric analysis. The introduction of analogous concepts of Ricci curvature on graphs brings new deep insights and tools to the study of graph theory. In [28, 29], Ollivier introduced the concept of coarse Ricci curvature for Markov chains on metric spaces, including graphs. In [23], Lin, Lu, and Yau introduced a modified version of Ollivier’s coarse Ricci curvature on graphs. This concept of modified curvature, which we call the Lin–Lu–Yau curvature, has been proved to be very useful to detect geometric and analytic properties of the underlying graphs [14, 13, 12, 25, 26, 27].
Explicit calculations of Lin–Lu–Yau curvature on graphs with local regularity restrictions have important applications in the study of locally correctable and some locally testable binary linear codes [20], and in bounding the diameter and eigenvalues of amply regular graphs [22, 19]. A $d$-regular graph $G$ with $n$ vertices is called an amply regular graph with parameters $(n,d,\alpha,\beta)$ if any two adjacent vertices have $\alpha$ common neighbors and any two vertices of distance $2$ have $\beta$ common neighbors. Amply regular graphs with diameter $2$ are strongly regular. For any two adjacent vertices, denote by $\kappa(x,y)$ the Lin–Lu–Yau curvature of the edge $xy$. In the works of [22, 19], it has been shown that for any two adjacent vertices $x,y$ of an amply regular graph $G$ with parameters $(n,d,\alpha,\beta)$, the following lower bound estimates hold:
  • If $\beta=\alpha>1$, then $\kappa(x,y)\geq\frac{2}{d}$.
  • If $\beta>\alpha\geq 1$, then $\kappa(x,y)\geq\frac{3}{d}$.
The Lin–Lu–Yau curvature of amply regular graphs with $\alpha>\beta\geq 2$ is rather mysterious. It is still open whether such graphs have positively lower bounded Lin–Lu–Yau curvature or not.
On the other hand, the Lin–Lu–Yau curvature has a natural upper bound. Achieving such an upper bound is equivalent to the existence of certain local perfect matchings. This is formulated precisely as below (see, e.g., [14, Proposition 2.7]).
Theorem 1.1.
Let $G=(V,E)$ be a $d$-regular graph. For any edge $xy\in E$, we have
$\kappa(x,y)\leq\frac{2+|\Delta_{xy}|}{d},$
where $\Delta_{xy}:=\Gamma(x)\cap\Gamma(y)$, $\Gamma(x):=\{z\in V|xz\in E\}$, and $\Gamma(y):=\{z\in V|yz\in E\}$. Moreover, the following two propositions are equivalent:
  • (a)
    $\kappa(x,y)=\frac{2+|\Delta_{xy}|}{d}$;
  • (b)
    there is a perfect matching between $N_{x}$ and $N_{y}$, where $N_{x}:=\Gamma(x)\backslash(\Delta_{xy}\cup\{y\})$ and $N_{y}:=\Gamma(y)\backslash(\Delta_{xy}\cup\{x\})$.
The crucial role of local matching conditions in the curvature calculation has been first observed in [3, 32], and further developed in [5, 15, 27]. For an amply regular graph $G$ with parameters $(n,d,\alpha,\beta)$, the above upper bound reads $\kappa(x,y)\leq\frac{2+\alpha}{d}$. A natural question is for which parameters this upper bound is achieved.
Bonini et al. [5] derived formulas for the Lin–Lu–Yau curvature of strongly regular graphs based on the parameters of graphs and the size of a maximal matching in the core neighborhood. They further proposed the following conjecture [5, Conjecture 1.7].
Conjecture 1.2.
Let $G=(V,E)$ be a strongly regular conference graph with parameters $(4\gamma+1,2\gamma,\gamma-1,\gamma)$, with $\gamma\geq 2$. Then the Lin–Lu–Yau curvature satisfies
$\kappa(x,y)=\frac{1}{2}+\frac{1}{2\gamma},\,\,\text{for all}\,\,xy\in E.$
Recall that a conference graph is a strongly regular graph with parameters $(4\gamma+1,2\gamma,\gamma-1,\gamma)$. Conference graphs have very interesting properties: While being highly symmetric, they are quasirandom [1, Theorem 9.3.1]. The existence of conference graphs are known for small values of $\gamma$ and for each $\gamma$ such that $14\gamma+$ 1 is a prime power (Paley graphs). Let $n=4\gamma+1$ be a prime power and $GF(n)$ be the unique finite field of order $n$. Recall that a Paley graph $P(n)$ of order $n$ is a graph on $n$ vertices with two vertices adjacent if their difference is a non-zero square in $GF(n)$. However, there are still many values of $\gamma$ for which the existence of conference graphs is open. Observe by Theorem 1.1 that the Lin–Lu–Yau curvature of a conference graph is no greater than $\frac{2+(\gamma-1)}{2\gamma}=\frac{1}{2}+\frac{1}{2\gamma}$. Therefore, Conjecture 1.2 claims the upper bound is achieved for conference graphs.
We stress that, in general, the Lin–Lu–Yau curvature of an amply regular graph is not necessarily uniquely determined by its parameters $(n,d,\alpha,\beta)$. For instance, the $4\times 4$ Rook’s graph and the Shrikhande graph are both strongly regular with parameters $(16,6,2,2)$, but Bonini et al. [5] calculated their Lin–Lu–Yau curvatures to be $\kappa=\frac{2}{3}$ and $\kappa=\frac{1}{3}$, respectively. In particular, the $4\times 4$ Rook’s graph takes the upper bound of the Lin–Lu–Yau curvature, while the Shrikhande graph does not.
Huang, Liu, and Xia showed [19, Proposition 3.6] that $\kappa(x,y)=\frac{2+\alpha}{d}$ for any amply regular graph with parameters $(n,d,\alpha,\beta)$ such that $d\leq 2\beta-\alpha-1$. However, the parameters of conference graphs violate the restriction $d\leq 2\beta-\alpha-1$. The lower curvature bound of amply regular graph with $\beta>\alpha\geq 1$ in [19], which has been mentioned earlier, applies to conference graphs and provides a positive lower bound $\frac{3}{d}$.
After finishing the first version of this paper, we notice from a very recent preprint [6] that the Lin–Lu–Yau curvature of the special case of Paley graphs $P(n)$ of even power order $n=p^{2m}$ has been confirmed to be $\frac{2+\alpha}{d}$ in an unpublished manuscript by Bonini, Liu, Semmens and Tran. Very interesting extensions to generalized Paley graphs have been carried out in [6].
Our main result in this paper confirms the Conjecture 1.2.
Theorem 1.3.
The Conjecture 1.2 holds true.
Indeed, our method depends only on the parameter relations and applies to more general classes of amply regular graphs. Below, we provide several sufficient parameter conditions.
Theorem 1.4.
Let $G=(V,E)$ be an amply regular graph with parameters $(n,d,\alpha,\beta)$. Suppose that one of the following conditions holds:
  • (1)
    $\beta=\alpha+1$, $n<3d-2\alpha$ and $d>\alpha+\sqrt{6\alpha+\frac{1}{4}}+\frac{3}{2}$;
  • (2)
    $\beta=\alpha+2$, $n<3d-2\alpha$ and $d\geq\alpha+\sqrt{10\alpha+\frac{41}{4}}+\frac{3}{2}$;
  • (3)
    $\beta>\alpha+2$, $n<3d-2\alpha$ and $\left|d-\frac{3}{2}\beta\right|\geq\frac{1}{2}\sqrt{4\alpha^{2}-3\beta^{2}+4\alpha+24\beta-20}$;
  • (4)
    $\beta\geq\frac{2\sqrt{3}}{3}\alpha+7$ and $n<3d-2\alpha$;
  • (5)
[…]
Then the Lin–Lu–Yau curvature satisfies
$\kappa(x,y)=\frac{2+\alpha}{d},\,\,\text{for all}\,\,xy\in E.$
By Theorem 1.1, Theorem 1.4 is equivalent to the following result in terms of local matchings.
Theorem 1.4.
Let $G=(V,E)$ be an amply regular graph with parameters $(n,d,\alpha,\beta)$. Suppose that one of the $5$ parameter conditions in Theorem 1.4 holds. Then there exists a perfect matching between $N_{x}$ and $N_{y}$ for all $xy\in E$.
Interestingly, the parameter conditions stated in Theorem 1.4 are fulfilled by many known strongly regular graphs, such as Paley graphs, McLaughlin graph and so on. We provide a list of more graphs (on at most $512$ vertices) satisfying one of the parameter conditions in Theorem 1.4 in the Appendix. In particular, by the condition $(1)$ of Theorem 1.4, Conjecture 1.2 holds true for $\gamma>6$. The condition $(5)$ applies to many graphs of the mysterious case of amply regular graphs with $\alpha>\beta$, see the Appendix.
Our proof provides a new combinatorial method for showing the existence of local perfect matchings. This new trick can also be employed to deal with many cases of amply regular graphs other than those in Theorem 1.4. For example, the restriction $n<3d-2\alpha$ is not always necessary. We refer to Remark 3.4 and Example 3.5 for more detailed discussions.
Applying Theorem 1.3 to Paley graphs, we obtain an interesting result of number theory. Theorem 1.3 tells that there exists a perfect matching between $N_{x}$ and $N_{y}$ for any adjacent vertices $x,y$ of a Paley graph $P(n)$ with $n>5$. This leads to the following consequence.
Corollary 1.5.
Let $n>5$ be a prime power congruent to $1$ modulo $4$. Consider two numbers $x,y\in GF(n)$ such that $x-y$ is a non-zero square in $GF(n)$. For any subset $S\subset GF(n)\setminus\{x,y\}$ with $|S|\geq\frac{3}{4}(n-1)$, there exist $w,z\in S$ such that $x-w,w-z,z-y$ are non-zero squares and $x-z,y-w$ are not squares in $GF(n)$.
In particular, this corollary applies to quadratic residues when $n>5$ is a Pythagorean prime, i.e., a prime congruent to $1$ modulo $4$.
Furthermore, we study another upper bound of the Lin–Lu–Yau curvature, which is given in terms of the Laplacaian eigenvalue and is an analogue to the Lichnerowicz Theorem in Riemannian geometry.
Theorem 1.6 (Discrete Lichnerowicz Theorem [28, 23]).
Let $G=(V,E)$ be a finite connected graph. Then the Lin–Lu–Yau curvature statisfies
$\min_{xy\in E}\kappa(x,y)\leq\lambda_{2},$
where $\lambda_{2}$ is the first nonzero eigenvalue of the normalized Laplacian $L$. Here the normalized Laplacian $L$ is defined as $L=I-D^{-\frac{1}{2}}AD^{-\frac{1}{2}}$, with $A$ being the adjacency matrix of the graph $G$ and $D$ being the diagonal matrix of vertex degrees.
Following [14], we say a graph $G$ is Lichnerowicz sharp if $\lambda_{2}=\min_{xy\in E}\kappa(x,y)$, i.e., the above upper bound is achieved. We observe the following interesting phenomena.
Theorem 1.7.
Let $G=(V,E)$ be a connected strongly regular graph with parameters $(n,d,\alpha,\beta)$ such that $\beta\geq\alpha$. If $G$ is Lichnerowicz sharp, i.e., $\min_{xy\in E}\kappa(x,y)=\lambda_{2}$, then
$\kappa(x,y)=\frac{2+\alpha}{d}.$
Indeed, all distance regular graphs with the two curvature upper bounds in Theorem 1.1 and Theorem 1.6 both achieved, i.e.,
$\min_{xy\in E}\kappa(x,y)=\lambda_{2}=\frac{2+\alpha}{d},$
have been fully classified in [14, Theorem 6.4]. Accordingly, we obtain the following classification of Licherowicz sharp strongly regular graphs with $\beta\geq\alpha$.
Corollary 1.8.
The connected Lichnerowicz sharp strongly regular graphs with parameters $(n,d,\alpha,\beta)$ such that $\beta\geq\alpha$ are precisely the following ones: The cocktail party graphs $CP(k),k\geq 2$, the Paley graph $P(9)$, the Johnson graphs $J(5,2)$ and $J(6,2)$ and the demi-cube $Q^{5}_{(2)}$ (also called Clebsch graph).
The rest of the our paper is organized as follows. In Section 2, we recall definitions and properties about the Lin–Lu–Yau curvature, perfect matching, and strongly/amply regular graphs. In Section 3, we prove Theorem 1.3, Theorem 1.4 and Corollary 1.5, and present possible ways to generalize our conclusion. In Section 4, we prove the Theorem 1.7 and Corollary 1.8.
2 Preliminaries
2.1 Lin–Lu–Yau curvature and local matchings
Definition 2.1 (Wasserstein distance).
Let $G=(V,E)$ be a locally finite graph, $\mu_{1}$ and $\mu_{2}$ be two probability measures on $G$. The Wasserstein distance $W_{1}(\mu_{1},\mu_{2})$ between $\mu_{1}$ and $\mu_{2}$ is defined as
$W_{1}(\mu_{1},\mu_{2})=\inf_{\pi}\sum_{y\in V}\sum_{x\in V}d(x,y)\pi(x,y),$
where $d(x,y)$ denotes the combinatorial distance between $x$ and $y$ in $G$, and the infimum is taken over all maps $\pi:V\times V\to[0,1]$ satisfying
$\mu_{1}(x)=\sum_{y\in V}\pi(x,y),\,\,\mu_{2}(y)=\sum_{x\in V}\pi(x,y).$
Such a map is called a transport plan.
We consider the following particular measure around a vertex $x\in V$:
$\mu_{x}^{p}(y)=\left\{\begin{array}[]{ll}p,&\hbox{if $y=x$;}\\ \frac{1-p}{\mathrm{deg}(x)},&\hbox{if $y\sim x$;}\\ 0,&\hbox{otherwise,}\end{array}\right.$
where $\mathrm{deg}(x):=\sum_{y\in V:y\sim x}1$ is the vertex degree of $x$.
Definition 2.2 ($p$-Ollivier curvature [28] and Lin–Lu–Yau curvature [23]).
Let $G=(V,E)$ be a locally finite graph. For any vertices $x,y\in V$, the $p$-Ollivier curvature $\kappa_{p}(x,y)$, $p\in[0,1]$, is defined as
$\kappa_{p}(x,y)=1-\frac{W_{1}(\mu_{x}^{p},\mu_{y}^{p})}{d(x,y)}.$
The Lin–Lu–Yau curvature $\kappa(x,y)$ is defined as
$\kappa(x,y)=\lim_{p\to 1}\frac{\kappa_{p}(x,y)}{1-p}.$
Notice that $\kappa_{1}(x,y)$ is always 00. Hence, the Lin–Lu–Yau curvature $\kappa(x,y)$ equals the negative of left derivative of the function $p\mapsto\kappa_{p}(x,y)$ at $p=1$.
Boure et al. [7] proved the following relation between $p$-Ollivier curvature and Lin–Lu–Yau curvature for an edge $xy\in E$ of a $d$-regular graph $G=(V,E)$:
$\kappa(x,y)=\frac{d+1}{d}\kappa_{\frac{1}{d+1}}(x,y).$ (2.1)
This relation leads in particular to the following expression of Lin–Lu–Yau curvature of a $d$-regular graph:
$\kappa(x,y)=\frac{1}{d}\left(d+1-\min_{\begin{subarray}{c}\phi:N_{x}\to N_{y}\\ \text{bijective}\end{subarray}}\sum_{v\in N_{x}}d(v,\phi(v))\right).$ (2.2)
Theorem 1.1 on the upper bound of Lin–Lu–Yau curvature and its relation to existence of a perfect matching between $N_{x}$ and $N_{y}$ follows then directly from (2.2). For more details, we refer to [14, Proposition 2.7].
Let us recall the definition of a perfect matching.
Definition 2.3.
[4, Section 16.1] Let $G=(V,E)$ be a locally finite simple connected graph. A set $M$ of pairwise non-adjacent edges is called a matching. Each vertex adjacent to an edge of $M$ is said to be covered by $M$. A matching $M$ is called a perfect matching if it covers every vertex of the graph.
Hall’s marriage Theorem is an fundamental tool for proving Conjecture 1.2.
Theorem 2.4.
[4, Theorem 16.4, Hall’s Marriage Theorem] Let $H=(V,E)$ be a bipartite graph with the bipartition $V=V_{1}\sqcup V_{2}$. Then $H$ has a perfect matching if and only if
$|V_{1}|=|V_{2}|\,\,\text{and}\,\,|\Gamma_{V_{2}}(S)|\geq|S|\,\,\text{for all}\,\,S\subseteq V_{1},$
holds, where $\Gamma_{V_{2}}(S):=\{v\in V_{2}|\,\,\text{there exists }\,w\in S\,\text{such that}\,\,vw\in E\}$.
2.2 Strongly/Amply regular graph
Definition 2.5 (Strongly/Amply regular graph [8]).
Let $G$ be a $d$-regular graph with $n$ vertices which is neither complete nor empty. Then $G$ is called call a strongly regular graph with parameters $(n,d,\alpha,\beta)$ if:
  • (i)
    Any two adjacent vertices have $\alpha$ common neighbors;
  • (ii)
    Any two non-adjacent vertices have $\beta$ common neighbors.
[…]
  • (i)
    Any two adjacent vertices have $\alpha$ common neighbors;
  • (ii)
    Any two vertices with distance $2$ have $\beta$ common neighbors.
The following theorem shows a relationship between the parameters of an amply regular graph.
Theorem 2.6.
Let $G=(V,E)$ be an amply regular graph with parameters $(n,d,\alpha,\beta)$. then
$d(d-\alpha-1)\leq(n-d-1)\beta.$ (2.3)
Moreover, the equality holds if and only if $G$ is strongly regular.
Proof.
Let $M$ be the number of pairs of ($(\{v_{1},v_{2}\},u)$, where $v_{1},v_{2},u\in V$ and $u$ is a common neighbour of $v_{1}$ and $v_{2}$. For any $u\in V$, there are $\binom{d}{2}$ pairs of such $\{v_{1},v_{2}\}$. Thus, $M=\binom{d}{2}n$. For any two adjacent vertices $v_{1}$ and $v_{2}$, there are $\alpha$ such $u$. For any two non-adjacent vertices $v_{1}$ and $v_{2}$, there are at either $\beta$ or 00 such $u$. Therefore,
$M\leq e\alpha+\left(\binom{n}{2}-e\right)\beta,$
where $e=|E|=\frac{nd}{2}$, and the inequality (2.3) follows. It’s clear that the equality holds if and only if $G$ is strongly regular. ∎
3 Proof of Theorem 1.3, Theorem 1.4 and Corollary 1.5
In this section, we prove Theorem 1.3, Theorem 1.4 and Corollary 1.5. First, we observe the following lemma which is useful in applying Hall’s marriage Theorem (Theorem 2.4).
Lemma 3.1.
Let $H=(V,E)$ be a bipartite graph with the bipartition $V=V_{1}\sqcup V_{2}$ such that $|V_{1}|=|V_{2}|$. Denote by $m:=|V_{1}|=|V_{2}|$. Suppose that
  • (i)
    $|\Gamma_{V_{2}}(S)|\geq|S|$ for any $S\subseteq V_{1}$ with $|S|\leq\frac{m+1}{2}$;
  • (ii)
    $|\Gamma_{V_{1}}(S)|\geq|S|$ for any $S\subseteq V_{2}$ with $|S|\leq\frac{m+1}{2}$.
Then, we have $|\Gamma_{V_{2}}(S)|\geq|S|$ for any $S\subset V_{1}$ and $|\Gamma_{V_{1}}(S)|\geq|S|$ for any $S\subset V_{2}$.
Proof.
Let $S\subseteq V_{1}$ be a subset of $V_{1}$ such that $|S|>\frac{m+1}{2}$. Pick a subset $S_{1}\subset S$ with $|S_{1}|=\left\lfloor\frac{m+1}{2}\right\rfloor$. It follows from the assumption (i) that
$|\Gamma_{V_{2}}(S)|\geq|\Gamma_{V_{2}}(S_{1})|\geq|S_{1}|=\left\lfloor\frac{m+1}{2}\right\rfloor.$
This implies that
$|V_{2}\setminus\Gamma_{V_{2}}(S)|=|V_{2}|-|\Gamma_{V_{2}}(S)|\leq m-\left\lfloor\frac{m+1}{2}\right\rfloor\leq\frac{m}{2}.$
Due to the assumption (ii), we have
$|\Gamma_{V_{1}}(V_{2}\setminus\Gamma_{V_{2}}(S))|\geq|V_{2}\setminus\Gamma_{V_{2}}(S)|.$
By definition, we have $\Gamma_{V_{1}}(V_{2}\setminus\Gamma_{V_{2}}(S))\cap S=\emptyset$. Thus, we estimate
$m=|V_{1}|\geq|\Gamma_{V_{1}}(V_{2}\setminus\Gamma_{V_{2}}(S))|+|S|\geq|V_{2}\setminus\Gamma_{V_{2}}(S)|+|S|=m-|\Gamma_{V_{2}}(S)|+|S|.$
Hence we have $|\Gamma_{V_{2}}(S)|\geq|S|$. This proves $|\Gamma_{V_{2}}(S)|\geq|S|$ for any $S\subseteq V_{1}$. By symmetry, we have $|\Gamma_{V_{1}}(S)|\geq|S|$ for any $S\subseteq V_{2}$. ∎
The following lemma is a direct consequence of Theorem 2.6.
Lemma 3.2.
Let $G=(V,E)$ be an amply regular graphs with parameters $(n,d,\alpha,\beta)$ such that $d\geq\alpha+2$. We have the following properties:
  • (i)
    If $n<3d-2\alpha$, then $d\leq 2\beta$.
  • (ii)
    If $2n<5d-3\alpha$, then $d\leq\frac{3}{2}\beta$.
Proof.
If $n<3d-2\alpha$, we derive from Theorem 2.6 that
$d(d-\alpha-1)\leq(n-d-1)\beta\leq 2(d-\alpha-1)\beta.$ (3.1)
Since $d\geq\alpha+2$, we deduce $d\leq 2\beta$. This proves $(i)$. $(ii)$ is deduced similarly. ∎
Let $x,y$ be two adjacent vertices of an amply regular graph $G=(V,E)$ with parameters $(n,d,\alpha,\beta)$. We decompose the vertex set $V$ into $6$ disjoint subsets as follows (depicted in Figure 1):
$V=\{x\}\cup\{y\}\cup\Delta_{xy}\cup N_{x}\cup N_{y}\cup P_{xy}.$
[…]
Figure 1: A schematic plot for the decomposition of $V$
The subsets above are defined by $\Delta_{xy}:=\Gamma(x)\cap\Gamma(y)$, $N_{x}:=\Gamma(x)\backslash(\Delta_{xy}\cup\{y\})$, $N_{y}:=\Gamma(y)\backslash(\Delta_{xy}\cup\{x\})$, and $P_{xy}:=V\backslash(\Gamma(x)\cup\Gamma(y))$, where $\Gamma(x):=\{z\in V|\,xz\in E\}$, and $\Gamma(y):=\{z\in V|\,yz\in E\}$. It is direct to check by definition that
$|\Delta_{xy}|=\alpha,\,|N_{x}|=|N_{y}|=d-\alpha-1,\,|P_{xy}|=n-2d+\alpha.$ (3.2)
Next, we prepare the following lemma.
Lemma 3.3.
Let $G=(V,E)$ be an amply regular graphs with parameters $(n,d,\alpha,\beta)$ and $x,y$ be two adjacent vertices. Suppose that a vertex $v\in N_{x}$ has $\ell$ neighbors in $N_{y}$. Then,
  • (i)
    $v$ has $\beta-1-\ell$ neighbors in $\Delta_{xy}$;
  • (ii)
    $v$ has $\alpha-\beta+1+\ell$ neighbors in $N_{x}$;
  • (ii)
    $v$ has $d-\alpha-1-\ell$ neighbors in $P_{xy}$.
Proof.
Since $y$ and $v$ have $\beta$ common neighbors, $v$ has $\beta-1-\ell$ neighbors in $\Delta_{xy}$. Since $x$ and $v$ have $\alpha$ common neighbors, $v$ has $\alpha-\beta+1+\ell$ neighbors in $N_{x}$. Since $v$ has $d-1$ neighbors except $x$, the number of neighbors of $v$ in $P_{xy}$ is $d-\alpha-1-\ell$. ∎
Now, we are prepared to prove Theorem 1.4.
Proof of Theorem 1.4.
Let $G=(V,E)$ be an amply regular graph satisfying one of the conditions (1)–(5), and $x,y\in V$ be two adjacent vertices. Recall from [19, Proposition 3.6] that $\kappa(x,y)=\frac{2+\alpha}{d}$ holds whenever $d\leq 2\beta-\alpha-1$. In the following proof, we assume that $d\geq 2\beta-\alpha$.
We claim that each of the conditions (1)–(5) yields $d>\alpha+2$. Indeed, it is clear that conditions (1) and (2) yield $d>\alpha+2$. Under the condition (3) or (4), $d\geq\beta>\alpha+2$. Under the condition (5), $d\geq 2\alpha+3-\frac{(\beta-1)^{2}}{\alpha}>\alpha+3$. Therefore, the claim holds. Recalling (3.2), we derive $|N_{x}|=|N_{y}|=d-\alpha-1>1$. That is, both $N_{x}$ and $N_{y}$ contain at least two vertices.
Recall from [22, Theorem 1.5] that $\kappa(x,y)=\frac{2+\alpha}{d}$ holds whenever $\alpha=0$ and $\beta\geq 2$. If $\alpha=0$, then each of the conditions (2)–(4) leads to $\beta\geq 2$ and hence $\kappa(x,y)=\frac{2}{d}$. Under the condition (1), $\alpha=0$ implies $\beta=1$ and $d>2$. However, by the condition $n<3d-2\alpha$ and Lemma 3.2, we have $d\leq 2\beta=2$, which is a contradiction. That is, under the condition (1), the case $\alpha=0$ is not allowed. Under the condition (5), it’s clear that the case $\alpha=0$ can not happen. In the following proof, we assume that $\alpha>0$.
By Theorem 1.1, we only need to prove that there is a perfect matching between $N_{x}$ and $N_{y}$. By Hall’s marriage Theorem (see Theorem 2.4), it remains to prove that, for any subset $S\subseteq N_{x}$, the number $|\Gamma_{N_{y}}(S)|$ of neighbors in $N_{y}$ is at least $|S|$. For simplicity, we denote $S:=\{v_{1},v_{2},\ldots,v_{b}\}$ with $b=|S|$ and $T:=\Gamma_{N_{y}}(S)$. We need to show the inequality
$|T|\geq b$ (3.3)
always holds. For each vertex $v_{i}\in S:=\{v_{1},v_{2},\ldots,v_{b}\}$, let $x_{i}$ be the number of neighbors of $v_{i}$ in $N_{y}$. By Lemma 3.3, the vertex $v_{i}$ has $d-\alpha-1-x_{i}$ neighbors in $P_{xy}$. If $|P_{xy}|=0$, then we have $x_{i}=d-\alpha-1=|N_{y}|$. Therefore, the inequality (3.3) clearly holds. In the following proof, we assume that $|P_{xy}|>0$, i.e., $n-2d+\alpha>0$.
By symmetry and Lemma 3.1, we only need to show the inequality (3.3) for $b\leq\frac{d-\alpha}{2}$.
We argue by contradiction. Suppose $|T|\leq b-1$.
Case 1: $b=1$.
If $b=1$, then $v_{1}$ has no neighbor in $N_{y}$, that is, $x_{1}=0$. Thus by Lemma 3.3, $v_{1}$ has $\beta-1$ neighbors in $\Delta_{xy}$ and $d-\alpha-1$ neighbors in $P_{xy}$. If condition (2) or (3) or (4) holds, then $\beta-1>\alpha=|\Delta_{xy}|$, which is a contradiction. If condition (5) holds, then $|P_{xy}|=n-2d+\alpha\leq\frac{1}{2}(d-\alpha-1)<d-\alpha-1$, which is a contradiction. If condition (1) holds, we have $\beta-1=\alpha$ and hence $v_{1}$ is adjacent to all vertices in $\Delta_{xy}$. Since $n<3d-2\alpha$, we have $|P_{xy}|=n-2d+\alpha\leq d-\alpha-1$. Thus, $|P_{xy}|=d-\alpha-1$ and $v_{1}$ is adjacent to all vertices in $P_{xy}$. Therefore, $v_{1}$ has in total $d-1$ neighbors in $\Delta_{xy}\cup P_{xy}$. Hence, $v_{1}$ has no neighbors in $N_{x}\cup N_{y}$.
We pick any vertex $v\in N_{x}\setminus\{v_{1}\}$, and suppose that $v$ has $p$ neighbors in $N_{y}$. Then $v$ has $\beta-p-1$ neighbors in $\Delta_{xy}$ and $d-\alpha-1-p$ neighbors in $P_{xy}$. Therefore, $v$ and $v_{1}$ have a total of $(\beta-1-p)+(d-\alpha-1-p)+1$ common neighbors. Since $v$ and $v_{1}$ are of distance $2$, the number of common neighbors of $v$ and $v_{1}$ equals to $\beta$. Thus,
$(\beta-1-p)+(d-\alpha-1-p)+1=\beta,$
that is,
$2p=d-\alpha-1,$ (3.4)
which implies that $d-\alpha-1$ is even.
Similarly, we pick any vertex $u\in N_{y}$, and suppose that $u$ has $q$ neighbors in $N_{x}$. Then $u$ has $\beta-q-1$ neighbors in $\Delta_{xy}$ and $d-\alpha-1-q$ neighbors in $P_{xy}$. Therefore, $u$ and $v_{1}$ have a total of $(\beta-q-1)+(d-\alpha-1-q)$ common neighbors. Since $u$ and $v_{1}$ are non-adjacent, we have $q<|N_{x}|=d-\alpha-1$. Thus, $u$ and $v_{1}$ have at least one common neighbor, and hence the number of common neighbors of $u$ and $v_{1}$ equals to $\beta$. Then,
$(\beta-q-1)+(d-\alpha-1-q)=\beta,$
that is,
$2\displaystyle 2q=d-\alpha-$ 2 . (3.5)
Thus, $d-\alpha-1$ is odd, which is a contradiction.
Case 2: $2\leq b\leq\frac{d-\alpha}{2}$.
Let $X=x_{1}+\cdots+x_{b}$, then $X$ is the number of edges between $S:=\{v_{1},\ldots,v_{b}\}$ and $T$.
Set $\Delta_{xy}:=\{\delta_{1},\ldots,\delta_{\alpha}\}$, and let $A_{k}$ be the number of neighbors of $\delta_{k}$ in $S$, where $1\leq k\leq\alpha$. Counting the number of edges between $S$ and $\Delta_{xy}$, we have
$\sum_{k=1}^{\alpha}A_{k}=\sum_{i=1}^{b}(\beta-x_{i}-1)=(\beta-1)b-X.$
Let $X_{ij}$ be the number of common neighbors of $v_{i}$ and $v_{j}$ in $\Delta_{xy}$, and denote by
$M_{1}:=\sum_{1\leq i<j\leq b}X_{ij}.$
Note that, for any $\delta_{k}\in\Delta_{xy}$, there are $\binom{A_{k}}{2}$ pairs of $v_{i},v_{j}$ so that $\delta_{k}$ is a common neighbor of them. We have
$M_{1}=\sum_{k=1}^{\alpha}\binom{A_{k}}{2}=\frac{1}{2}\left[\sum_{k=1}^{\alpha}A_{k}^{2}-\sum_{k=1}^{\alpha}A_{k}\right].$
By the Cauchy inequality, we derive
$M_{1}$ $\geq\frac{1}{2}\left[\frac{1}{\alpha}\left(\sum_{k=1}^{\alpha}A_{k}\right)^{2}-\sum_{k=1}^{\alpha}A_{k}\right]$
$=\frac{1}{2}\left[\frac{\left((\beta-1)b-X\right)^{2}}{\alpha}-\left((\beta-1)b-X\right)\right].$ (3.6)
Set $P_{xy}:=\{\epsilon_{1},\ldots,\epsilon_{n-2d+\alpha}\}$, and let $B_{k}$ be the number of neighbors of $\epsilon_{k}$ in $S$, where $1\leq k\leq n-2d+\alpha$. Similarly, we have
$\sum_{k=1}^{n-2d+\alpha}B_{k}=\sum_{i=1}^{b}(d-\alpha-1-x_{i})=(d-\alpha-1)b-X.$
Let $Y_{ij}$ be the number of common neighbors of $v_{i}$ and $v_{j}$ in $P_{xy}$, and denote by
$M_{2}:=\sum_{1\leq i<j\leq b}Y_{ij}.$
Since $n-2d+\alpha\leq d-\alpha-1$ under each condition $(1)-(5)$, it follows by Cauchy inequality that
$M_{2}=\sum_{k=1}^{n-2d+\alpha}\binom{B_{k}}{2}$ $\geq\frac{1}{2}\left[\frac{1}{n-2d+\alpha}\left(\sum_{k=1}^{n-2d+\alpha}B_{k}\right)^{2}-\sum_{k=1}^{n-2d+\alpha}B_{k}\right]$
$=\frac{1}{2}\left[\frac{\left((d-\alpha-1)b-X\right)^{2}}{n-2d+\alpha}-\left((d-\alpha-1)b-X\right)\right]$ (3.7)
$\geq\frac{1}{2}\left[\frac{\left((d-\alpha-1)b-X\right)^{2}}{d-\alpha-1}-\left((d-\alpha-1)b-X\right)\right].$ (3.8)
Set $T:=\{u_{1},\ldots,u_{|T|}\}$, and let $C_{k}$ be the number of neighbors of $u_{k}$ in $S$, where $1\leq k\leq|T|$. Similarly, we have
$\sum_{k=1}^{|T|}C_{k}=X.$
Let $Z_{ij}$ be the number of common neighbors of $v_{i}$ and $v_{j}$ in $T$, and denote by
$M_{3}:=\sum_{1\leq i<j\leq b}Z_{ij}.$
Similarly, we have
$M_{3}=\sum_{k=1}^{|T|}\binom{C_{k}}{2}=\frac{1}{2}\left[\sum_{k=1}^{|T|}C_{k}^{2}-\sum_{k=1}^{|T|}C_{k}\right].$
By the assumption that $|T|\leq b-1$, we derive
$M_{3}\geq\frac{1}{2}\left[\frac{1}{|T|}\left(\sum_{k=1}^{|T|}C_{k}\right)^{2}-\sum_{k=1}^{|T|}C_{k}\right]\geq\frac{1}{2}\left[\frac{1}{b-1}X^{2}-X\right].$ (3.9)
Set $N_{x}=\{v_{1},\ldots,v_{d-\alpha-1}\}$, and let $D_{k}$ be the number of neighbors of $v_{k}$ in $S$, where $11\leq k\leq d-\alpha-$ 1. Counting the number of ordered pairs $(v_{k},v_{j})$ with $v_{k}\in N_{x}$, $v_{j}\in S$ and $v_{k}v_{j}\in E$, we have
$\sum_{k=1}^{d-\alpha-1}D_{k}=\sum_{i=1}^{b}(\alpha-\beta+1+x_{i})=(\alpha-\beta+1)b+X.$
Let $W_{ij}$ be the number of common neighbors of $v_{i}$ and $v_{j}$ in $N_{x}$, and denote by
$M_{4}:=\sum_{1\leq i<j\leq b}W_{ij}.$
Similarly, we have
$M_{4}=\sum_{k=1}^{d-\alpha-1}\binom{D_{k}}{2}$ $\geq\frac{1}{2}\left[\frac{1}{d-\alpha-1}\left(\sum_{k=1}^{d-\alpha-1}D_{k}\right)^{2}-\sum_{k=1}^{d-\alpha-1}D_{k}\right]$
$=\frac{1}{2}\left[\frac{((\alpha-\beta+1)b+X)^{2}}{d-\alpha-1}-((\alpha-\beta+1)b+X)\right].$ (3.10)
Since $v_{i}$ and $v_{j}$ have either $\alpha$ or $\beta$ common neighbors, we have
$M_{1}+M_{2}+M_{3}+M_{4}+\binom{b}{2}$ $=\sum_{1\leq i<j\leq b}(X_{ij}+Y_{ij}+Z_{ij}+W_{ij}+1)$
$\leq\binom{b}{2}\max\{\alpha,\beta\}.$ (3.11)
We divide the remaining argument of Case 2 into four subcases.
(i) Suppose that the condition (1) holds. Inserting (3), (3.8), (3.9), (3) into (3) yields
$\left(\frac{1}{\alpha}+\frac{2}{d-\alpha-1}+\frac{1}{b-1}\right)X^{2}-4bX+(b^{2}-b)(d-\alpha-1)\leq 0.$ (3.12)
Since the coefficient of $X^{2}$ is positive, the discriminant of the above quadratic polynomial has to be non-negative. Therefore, we derive
$4b\geq$ $\left(\frac{1}{\alpha}+\frac{2}{d-\alpha-1}+\frac{1}{b-1}\right)(b-1)(d-\alpha-1)$
$=$ $2(b-1)+(d-\alpha-1)+\frac{(b-1)(d-\alpha-1)}{\alpha}.$ (3.13)
That is,
$d-\alpha-3-\frac{d-\alpha-1}{\alpha}\leq\left(2-\frac{d-\alpha-1}{\alpha}\right)b.$
Since $d\leq 2\beta\leq 3\alpha+1$, it follows that $2-\frac{d-\alpha-1}{\alpha}\geq 0$. By the assumption that $b\leq\frac{d-\alpha}{2}$, we have
$d-\alpha-3-\frac{d-\alpha-1}{\alpha}\leq\left(2-\frac{d-\alpha-1}{\alpha}\right)\frac{d-\alpha}{2}.$
Considering the above inequality as a quadratic inequality with respect to $d$, we deduce
$d\leq\alpha+\sqrt{6\alpha+\frac{1}{4}}+\frac{3}{2},$
which is a contradiction.
(ii) Suppose that the condition (2) holds. Then we derive from (3) that
$M_{1}\geq\frac{1}{2}\left[\frac{\big{(}(\alpha+1)b-X\big{)}^{2}}{\alpha+1}-\big{(}(\alpha+1)b-X\big{)}\right].$ (3.14)
Inserting (3.14), (3.8), (3.9), and (3) into (3), we have
$\left(\frac{1}{\alpha+1}+\frac{2}{d-\alpha-1}+\frac{1}{b-1}\right)X^{2}$ $-\left(4b+\frac{2b}{d-\alpha-1}\right)X$
$+\frac{b^{2}}{d-\alpha-1}+b+(b^{2}-b)(d-\alpha-1)\leq 0.$
Since the discriminant of the above quadratic polynomial is non-negative, we deduce that
$b\left(2+\frac{1}{d-\alpha-1}\right)^{2}$ $\geq\left[\frac{1}{\alpha+1}+\frac{2}{d-\alpha-1}+\frac{1}{b-1}\right]\left[\frac{b}{d-\alpha-1}+1+(b-1)(d-\alpha-1)\right]$
$>2(b-1)+d-\alpha-1+\frac{b}{(d-\alpha-1)^{2}}+\frac{2}{d-\alpha-1}+\frac{(b-1)(d-\alpha-1)}{\alpha+1}.$
That is,
$0>d-\alpha-2+\frac{2}{d-\alpha-1}-\frac{d}{\alpha+1}-\left(3+\frac{4}{d-\alpha-1}-\frac{d}{\alpha+1}\right)b.$ (3.15)
Since $d\leq 2\beta$, we obtain
$3+\frac{4}{d-\alpha-1}-\frac{d}{\alpha+1}>3-\frac{d}{\alpha+1}\geq 3-\frac{2\beta}{\alpha+1}\geq 0,$
and hence
00 0 $>d-\alpha-2+\frac{2}{d-\alpha-1}-\frac{d}{\alpha+1}-\left(3+\frac{4}{d-\alpha-1}-\frac{d}{\alpha+1}\right)\frac{d-\alpha}{2}$
$=d-\alpha-4-\frac{3(d-\alpha)}{2}+\frac{d(d-\alpha-2)}{2(\alpha+1)}.$ (3.16)
Considering the above inequality as a quadratic inequality with respect to $d$, we deduce
$d<\alpha+\sqrt{10\alpha+\frac{41}{4}}+\frac{3}{2},$
which is a contradiction.
(iii) Suppose that either condition (3) or (4) holds. Inserting (3), (3.8), (3.9) and $M_{4}\geq 0$ into the inequality (3) leads to
$\left(\frac{1}{\alpha}+\frac{1}{d-\alpha-1}+\frac{1}{b-1}\right)X^{2}$ $-\left(\frac{2b(\beta-1)}{\alpha}+2b-1\right)X$
$+\frac{b^{2}(\beta-1)^{2}}{\alpha}-(\beta-1)b+(b^{2}-b)(d-\alpha-\beta)\leq 0.$
Using $\frac{2b(\beta-1)}{\alpha}+2b-1<\frac{2b(\beta-1)}{\alpha}+2b$, we derive from the non-negativity of the discriminant of the above quadratic polynomial that
$b\left(\frac{\beta-1}{\alpha}+1\right)^{2}>\left[\frac{1}{\alpha}+\frac{1}{d-\alpha-1}+\frac{1}{b-1}\right]\left[\frac{b(\beta-1)^{2}}{\alpha}-(\beta-1)+(b-1)(d-\alpha-\beta)\right].$
That is,
$\frac{2b(\beta-1)}{\alpha}+2b>$ $1+\frac{bd-b\beta-d+1}{\alpha}$
$+\left[\frac{1}{d-\alpha-1}+\frac{1}{b-1}\right]\left[\frac{b(\beta-1)^{2}}{\alpha}-(\beta-1)+(b-1)(d-\alpha-\beta)\right].$
Since
$\frac{b(\beta-1)^{2}}{\alpha}-(\beta-1)+(b-1)(d-\alpha-\beta)=(b-1)(d-\alpha-1)+b(\beta-1)\left(\frac{\beta-1}{\alpha}-1\right),$
we derive
$\frac{2b(\beta-1)}{\alpha}+b$
$>$ $\frac{bd-b\beta-d+1}{\alpha}+d-\alpha-1+\left(\frac{1}{d-\alpha-1}+\frac{1}{b-1}\right)b(\beta-1)\left(\frac{\beta-1}{\alpha}-1\right)$
$>$ $\frac{bd-b\beta-d+1}{\alpha}+d-\alpha-1+(\beta-1)\left(\frac{\beta-1}{\alpha}-1\right)+\frac{b(\beta-1)}{d-\alpha-1}\left(\frac{\beta-1}{\alpha}-1\right).$
That is,
$\frac{-d+1}{\alpha}+d-\alpha-1$ $+(\beta-1)\left(\frac{\beta-1}{\alpha}-1\right)$
$<\left[\frac{2(\beta-1)}{\alpha}+1-\frac{d-\beta}{\alpha}-\frac{\beta-1}{d-\alpha-1}\left(\frac{\beta-1}{\alpha}-1\right)\right]b.$
Since $2\beta-\alpha\leq d\leq 2\beta\leq\alpha+\frac{5}{2}\beta-\frac{3}{2}$, we have
$\frac{2(\beta-1)}{\alpha}+1-\frac{d-\beta}{\alpha}-\frac{(\beta-1)}{d-\alpha-1}\left(\frac{\beta-1}{\alpha}-1\right)$
$\geq$ $\frac{2(\beta-1)}{\alpha}-\frac{3(\beta-1)}{2\alpha}-\frac{\beta-1}{2(\beta-\alpha-1)}\left(\frac{\beta-1}{\alpha}-1\right)=0.$
Therefore, we obtain
$\frac{-d+1}{\alpha}+d-\alpha-1+(\beta-1)\left(\frac{\beta-1}{\alpha}-1\right)$
$<$ $\left[\frac{2(\beta-1)}{\alpha}+1-\frac{d-\beta}{\alpha}-\frac{(\beta-1)}{d-\alpha-1}\left(\frac{\beta-1}{\alpha}-1\right)\right]\frac{d-\alpha}{2}$
$<$ $\left[\frac{2(\beta-1)}{\alpha}+1-\frac{d-\beta}{\alpha}\right]\frac{d-\alpha}{2}-\frac{(\beta-1)}{2}\left(\frac{\beta-1}{\alpha}-1\right).$
That is,
$d^{2}-3\beta d-\alpha^{2}-\alpha+3\beta^{2}-6\beta+5<0.$ (3.17)
Since the discriminant of the above quadratic polynomial is non-negative, we deduce that
$9\beta^{2}\geq 4(-\alpha^{2}-\alpha+3\beta^{2}-6\beta+5).$
Thus, we arrive at
$\beta\leq\sqrt{\frac{1}{3}(2\alpha+1)^{2}+9}+4<\frac{2\sqrt{3}}{3}\alpha+7.$
By (3.17), we have
$\left|d-\frac{3}{2}\beta\right|<\frac{1}{2}\sqrt{4\alpha^{2}-3\beta^{2}+4\alpha+24\beta-20}).$
Therefore, neither condition (3) nor (4) holds, which is a contradiction.
(iv) Suppose that the condition (5) holds. Then we have $n-2d+\alpha\leq\frac{1}{2}(d-\alpha-1)$. It follows from (3.7) that
$M_{2}\geq\frac{1}{2}\left[\frac{2\left((d-\alpha-1)b-X\right)^{2}}{d-\alpha-1}-\left((d-\alpha-1)b-X\right)\right].$ (3.18)
Inserting (3), (3.18), (3.9) and (3) into the inequality (3), we have
$\left(\frac{1}{\alpha}+\frac{3}{d-\alpha-1}+\frac{1}{b-1}\right)X^{2}-\left(\frac{2b(\beta-1)}{\alpha}+4b-\frac{2(\alpha-\beta+1)b}{d-\alpha-1}\right)X$
$\hskip 19.91684pt+\frac{b^{2}(\beta-1)^{2}}{\alpha}+2(d-\alpha-1)b^{2}+\frac{(\alpha-\beta+1)^{2}b^{2}}{d-\alpha-1}-(d-\alpha)b-(\alpha-1)b^{2}\leq 0.$ (3.19)
Since the discriminant of the above quadratic polynomial is non-negative, we deduce
$b\left(\frac{\beta-1}{\alpha}+2-\frac{\alpha-\beta+1}{d-\alpha-1}\right)^{2}\geq\left(\frac{1}{\alpha}+\frac{3}{d-\alpha-1}+\frac{1}{b-1}\right)\times$
$\hskip 19.91684pt\left(\frac{b(\beta-1)^{2}}{\alpha}+2(d-\alpha-1)b+\frac{(\alpha-\beta+1)^{2}b}{d-\alpha-1}-(d-\alpha)-(\alpha-1)b\right).$ (3.20)
It is clear that
$\frac{1}{\alpha}\left(\frac{b(\beta-1)^{2}}{\alpha}+2(d-\alpha-1)b+\frac{(\alpha-\beta+1)^{2}b}{d-\alpha-1}-(d-\alpha)-(\alpha-1)b\right)$
$>$ $\frac{b(\beta-1)^{2}}{\alpha^{2}}+\frac{2b(d-\alpha-1)}{\alpha}-\frac{d-\alpha+\alpha b-b}{\alpha}.$ (3.21)
Since $d\geq 2\alpha+3-\frac{(\beta-1)^{2}}{\alpha}$, we have
$d-\alpha-1\geq\alpha+2-\frac{(\beta-1)^{2}}{\alpha}>\alpha-\beta+3\geq 3.$ (3.22)
Thus, we estimate
$\frac{3}{d-\alpha-1}\left(\frac{b(\beta-1)^{2}}{\alpha}+2(d-\alpha-1)b+\frac{(\alpha-\beta+1)^{2}b}{d-\alpha-1}-(d-\alpha)-(\alpha-1)b\right)$
$=$ $\frac{3b(\beta-1)^{2}}{(d-\alpha-1)\alpha}+6b+\frac{3(\alpha-\beta+1)^{2}b}{(d-\alpha-1)^{2}}-3-\frac{3}{d-\alpha-1}-\frac{3(\alpha-1)b}{d-\alpha-1}$
$>$ $\frac{3b(\beta-1)^{2}}{(d-\alpha-1)\alpha}+6b+\frac{(\alpha-\beta+1)^{2}b}{(d-\alpha-1)^{2}}-4-\frac{3(\alpha-1)b}{d-\alpha-1}.$ (3.23)
By the condition that $d\geq 2\alpha+3-\frac{(\beta-1)^{2}}{\alpha}$, we derive
$\frac{1}{b-1}\left(\frac{b(\beta-1)^{2}}{\alpha}+2(d-\alpha-1)b+\frac{(\alpha-\beta+1)^{2}b}{d-\alpha-1}-(d-\alpha)-(\alpha-1)b\right)$
$=$ $\frac{(\beta-1)^{2}}{\alpha}+2(d-\alpha-1)+\frac{(\alpha-\beta+1)^{2}}{d-\alpha-1}-(\alpha-1)$
$\hskip 39.83368pt+\frac{1}{b-1}\left(\frac{(\beta-1)^{2}}{\alpha}+2(d-\alpha-1)+\frac{(\alpha-\beta+1)^{2}}{d-\alpha-1}-(d-1)\right)$
$>$ $\frac{(\beta-1)^{2}}{\alpha}+2(d-\alpha-1)+\frac{(\alpha-\beta+1)^{2}}{d-\alpha-1}-(\alpha-1)$
$>$ $\frac{(\beta-1)^{2}}{\alpha}+2(d-\alpha-1)+\frac{2(\alpha-\beta+1)^{2}}{\alpha}-(\alpha-1),$ (3.24)
where the last inequality above follows from Lemma 3.2. Inserting (3.21),(3),(3) into (3.20) yields
$-\frac{d}{\alpha}+\frac{(\beta-1)^{2}}{\alpha}+2(d-\alpha-1)+\frac{2(\alpha-\beta+1)^{2}}{\alpha}-\alpha-2$
$<$ $\left(\frac{4\beta-2d-3}{\alpha}-\frac{\alpha-\beta+4}{d-\alpha-1}+\frac{(\beta-1)(\alpha-\beta+1)}{\alpha(d-\alpha-1)}+1\right)b.$ (3.25)
By Lemma 3.2, we have $\frac{3}{2}\beta\geq d$. By (3.22), we have $d-\alpha-1\geq\alpha-\beta+4$. Thus, the coefficient of $b$ satisfies
$\frac{4\beta-2d-3}{\alpha}-\frac{\alpha-\beta+4}{d-\alpha-1}+\frac{(\beta-1)(\alpha-\beta+1)}{\alpha(d-\alpha-1)}+1\geq\frac{\beta-3}{\alpha}-1+0+1\geq 0.$
It follows from (3) that
$-\frac{d}{\alpha}+\frac{(\beta-1)^{2}}{\alpha}+2(d-\alpha-1)+\frac{2(\alpha-\beta+1)^{2}}{\alpha}-\alpha-2$
$<$ $\left(\frac{4\beta-2d-3}{\alpha}-\frac{\alpha-\beta+4}{d-\alpha-1}+\frac{(\beta-1)(\alpha-\beta+1)}{\alpha(d-\alpha-1)}+1\right)\frac{d-\alpha}{2}$
$\leq$ $\frac{(4\beta-2d-3)(d-\alpha)}{2\alpha}-\frac{\alpha-\beta+4}{2}+\frac{(\beta-1)(\alpha-\beta+1)}{2\alpha}+\frac{(\beta-1)(\alpha-\beta+1)}{2\alpha(d-\alpha-1)}+\frac{d-\alpha}{2}$
$<$ $\frac{(4\beta-2d-3)(d-\alpha)}{2\alpha}-\frac{\alpha-\beta+4}{2}+\frac{(\beta-1)(\alpha-\beta+1)}{2\alpha}+\frac{1}{2}+\frac{d-\alpha}{2}.$ (3.26)
That is,
$02d^{2}+(\alpha-4\beta+1)d-6\alpha\beta+\alpha+7\beta^{2}-14\beta+7<$ 0 .
Since $06\beta-d-1>$ 0, we have
$\alpha>\frac{2d^{2}-4\beta d+7\beta^{2}+d-14\beta+7}{6\beta-d-1},$
which is a contradiction.
Now, we have completed Case 2 and established Theorem 1.4.∎
Remark 3.4.
In addition to the cases in Theorem 1.4, our method can deal with many other cases. Here, we comment on some possible ways to extend our conclusion.
  • (i)
    The condition $n<3d-2\alpha$ is not always necessary to apply our method. In particular, one can employ (3.7) instead of (3.8).
  • (ii)
    Suppose that there are $e$ edges within $S:=\{v_{1},\ldots,v_{b}\}$ and $f$ edges between $S$ and $N_{x}\backslash S$. Note that $\sum_{k=1}^{b}D_{k}=2e$ and $\sum_{k=b+1}^{d-\alpha-1}D_{k}=f$. The inequality (3) can be improved as
    $M_{4}=\sum_{k=1}^{d-\alpha-1}\binom{D_{k}}{2}\geq\frac{1}{2}\biggl{(}\frac{(2e)^{2}}{b}+\frac{f^{2}}{d-\alpha-1-b}-\sum_{k=1}^{d-\alpha-1}D_{k}\biggr{)},$
    and the inequality (3) can be replaced by
    $M_{1}+M_{2}+M_{3}+M_{4}+\binom{b}{2}=\alpha e+\left(\binom{b}{2}-e\right)\beta.$
    These modifications might improve our conclusion to some extent.
To illustrate the above remark, we show that the Lin–Lu–Yau curvature of a particular strongly regular graph with $n\geq 3d-2\alpha$ achieves the upper bound $\frac{2+\alpha}{d}$.
Example 3.5.
Consider a strongly regular graph $G$ with parameters (324, 152, 70, 72) derived from regular symmetric Hadamard matrices with constant diagonal (RSHCD), which is constructed in [11, Section 5.3.2] based on [18, Lemma 11]. Note that the condition $n<3d-2\alpha$ does not hold. Inserting the parameters into (3), (3.7), (3.9), (3), and (3), we have
$\left(\frac{107}{5670}+\frac{1}{2(b-1)}\right)X^{2}-\frac{5462b}{2835}X+\frac{104791b^{2}}{2835}-40b\leq 0,$
where $412\leq b\leq\frac{d-\alpha}{2}=$ 41. Considering the discriminant, we have
$\left(\frac{5462b}{2835}\right)^{2}\geq 4\left(\frac{107}{5670}+\frac{1}{b-1}\right)\left(\frac{104791b^{2}}{2835}-40b\right).$
It is direct to check that the inequality above can not hold for $2\leq b\leq 41$. Therefore, the Lin–Lu–Yau curvature of $G$ is
$\frac{2+\alpha}{d}=\frac{9}{19}.$
Now, let’s prove Theorem 1.3. Recall that, in the setting of Theorem 1.3, we have $(n,d,\alpha,\beta)=(4\gamma+1,2\gamma,\gamma-1,\gamma)$. By the condition $(1)$ of Theorem 1.4, we only need to verify the cases that $\gamma$ is equal to $2$, $3$, $4$, $5$, or $6$.
Proof of Theorem 1.3.
We use the framework of the proof of Theorem 1.4. It is sufficient to derive contradiction assuming $2\leq b\leq\frac{d-\alpha}{2}=\frac{\gamma+1}{2}$, which implies $\gamma\geq 3$. By (3), we deduce directly that $b>\gamma-3$.
Suppose that $\gamma=3$, then $b=2$. That is, $\{v_{1},v_{2}\}$ have only $1$ neighbor (denoted by $u_{1}$) in $N_{y}$. We can assume both $v_{1}$ and $v_{2}$ are adjacent to $u_{1}$. Since, otherwise, one of them has no neighbors in $N_{y}$. This reduces to the Case $b=1$, which leads to a contradiction as in the proof of Theorem 1.4. Then, by Lemma 3.3, both $v_{1}$ and $v_{2}$ have $2$ neighbors in $P_{xy}$. Since $P_{xy}$ only has $3$ elements, $v_{1}$ and $v_{2}$ have at least $1$ common neighbor in $P_{xy}$. Since $x$ and $u_{1}$ are also common neighbors of $v_{1}$ and $v_{2}$, it follows that $v_{1}$ and $v_{2}$ have at least $3$ common neighbors. Thus, $v_{1}$ is non-adjacent to $v_{2}$. Since both $v_{1}$ and $v_{2}$ have a neighbor in $N_{x}$, they are both adjacent to $v_{3}:=N_{x}\setminus\{v_{1},v_{2}\}$. Hence, $v_{1}$ and $v_{2}$ have $4$ common neighbors, which contradicts to $\beta=\gamma=3$.
Suppose that $\gamma=4$, then $b=2$. That is, $\{v_{1},v_{2}\}$ have only $1$ neighbor in $N_{y}$. Clearly, $X=2$. By (3), (3.8) and (3.9), we have $M_{1}\geq 1$, $M_{2}\geq 2$ and $M_{3}\geq 1$. Hence,
$M_{1}+M_{2}+M_{3}+M_{4}+{\binom{b}{2}}>{\binom{b}{2}}\gamma,$
which is a contradiction.
Suppose that $\gamma=5$, then $b=3$. Substituting $\gamma=5$ and $b=3$ into the inequality (3.12), we have
$30-12X+\frac{23X^{2}}{20}\leq 0$
It follows that $X=5$ or $X=6$.
In case that $X=5$, we derive from (3), (3.8) and (3.9) that $M_{1}\geq 3$, $M_{2}\geq 5$ and $M_{3}\geq 4$ respectively. We claim that $M_{4}\geq 1$. Set $N_{x}=\{v_{1},\ldots,v_{5}\}$. Denote by $e$ the number of edges within $\{v_{1},\ldots,v_{5}\}$ and $f$ edges between $\{v_{1},\ldots,v_{5}\}$ and $N_{x}\backslash\{v_{1},\ldots,v_{5}\}$. Observe that $X=2e+f$ by Lemma 3.3 (ii). Thus, the number of edges within $\{v_{1},v_{2},v_{3}\}$ is at most $2$. If there are 2 edges within $\{v_{1},v_{2},v_{3}\}$, then $M_{4}\geq 1$, as claimed. If the number of edges within $\{v_{1},v_{2},v_{3}\}$ is at most $1$, then there are at least $3$ edges between $\{v_{1},v_{2},v_{3}\}$ and $\{v_{4},v_{5}\}$, which implies that either $v_{4}$ or $v_{5}$ is adjacent to two vertices of $\{v_{1},v_{2},v_{3}\}$. That is, we have $M_{4}\geq 1$ as claimed.
In case that $X=6$, we derive from (3), (3.8), (3.9) and (3) that $M_{1}\geq 2$, $M_{2}\geq 4$, $M_{3}\geq 6$ and $M_{4}\geq 1$ respectively.
Therefore, in either of the two cases $X=5$ and $x=6$, we have
$M_{1}+M_{2}+M_{3}+M_{4}+{\binom{b}{2}}>{\binom{b}{2}}\gamma,$
which is a contradiction.
Suppose that $\gamma=6$, then $b\leq\frac{7}{2}$. This contradicts to is contradictory to $b>\gamma-3$.
We have completed the proof. ∎
To conclude this section, we prove the number theoretic result Corollary 1.5.
Proof of Corollary 1.5.
Consider the corresponding Paley graph $P(n)$. By Theorem 1.1 and Theorem 1.3, there is a perfect matching $\mathcal{M}$ between $N_{x}$ and $N_{y}$. Since $|S|\geq\frac{3}{4}(n-1)$, we have $\left|\left(GF(n)\setminus\{x,y\}\right)\setminus S\right|\leq\frac{1}{4}(n-1)-1$. Recall a Paley graph $P(n)$ with $n=4\gamma+1$ is a strongly regular graph with parameters $(4\gamma+1,2\gamma,\gamma-1,\gamma)$. Therefore, we have $|N_{x}|=\gamma=\frac{1}{4}(n-1)$. That is, we have $\left|\left(GF(n)\setminus\{x,y\}\right)\setminus S\right|\leq|N_{x}|-1$. Then it is clear that $GF(n)\backslash S$ is not a vertex cover of $\mathcal{M}$. Suppose that $wz\in\mathcal{M}$ is not covered by $GF(n)\setminus S$, then the result follows. ∎
4 Proof of Theorem 1.7 and Corollary 1.8
In this section, we prove Theorem 1.7 and Corollary 1.8. We first recall the following facts about eigenvalues of strongly regular graphs, see, e.g., [17, Page 220].
Theorem 4.1.
Let $G=(V,E)$ be a strongly regular graph with parameters $(n,d,\alpha,\beta)$. Then the Laplacian eigenvalues of $G$ are
$\lambda_{1}=0,$
$\lambda_{2}=1-\frac{(\alpha-\beta)+\sqrt{(\alpha-\beta)^{2}+4(d-\beta)}}{2d},$
$\lambda_{3}=1-\frac{(\alpha-\beta)-\sqrt{(\alpha-\beta)^{2}+4(d-\beta)}}{2d}.$
The corresponding multiplicities are
$m_{1}=1,$
$m_{2}=\frac{1}{2}\biggl{(}(n-1)-\frac{2d+(n-1)(\alpha-\beta)}{\sqrt{(\alpha-\beta)^{2}+4(d-\beta)}}\biggr{)},$
$m_{3}=\frac{1}{2}\biggl{(}(n-1)+\frac{2d+(n-1)(\alpha-\beta)}{\sqrt{(\alpha-\beta)^{2}+4(d-\beta)}}\biggr{)}.$
By Theorem 1.1, Theorem 1.7 is a direct consequence of the following result.
Theorem 4.2.
Let $G=(V,E)$ be a strongly regular graph with parameters $(n,d,\alpha,\beta)$ such that $\beta\geq\alpha$. If $\lambda_{2}\leq\frac{2+\alpha}{d}$, then the equality holds, i.e., $\lambda_{2}=\frac{2+\alpha}{d}$.
Proof.
Since $\lambda_{2}\leq\frac{2+\alpha}{d}$, we have by Theorem 4.1 that
$2d-3\alpha+\beta-4\leq\sqrt{(\alpha-\beta)^{2}+4(d-\beta)}.$ (4.1)
Since $G$ is connected and not complete, we have $d\geq\alpha+2$. Therefore, we derive
$2d-3\alpha+\beta-4\geq 2(\alpha+2)-3\alpha+\beta-4\geq 0$
In the second inequality above, we use our assumption $\beta\geq\alpha$. It follows that we can square both sides of the inequality (4.1). Hence, we have
$(2d-3\alpha+\beta-4)^{2}\leq(\alpha-\beta)^{2}+4(d-\beta).$
The above inequality can be rearranged as
$4(d-\alpha-1)(d-2\alpha+\beta-4)\leq 0$
Since $d\geq\alpha+2$, we deduce
$d\leq 2\alpha-\beta+4.$ (4.2)
On one hand, since $\beta\geq\alpha$, we have $d\leq\alpha+4$. That is, $\alpha+2\leq d\leq\alpha+4$. On the other hand, since $d\geq\beta$, we have $\beta\leq\alpha+2$.
Observe that the multiplicity numbers $m_{2}$ and $m_{3}$ in Theorem 4.1 must be positive integers. Hence, $(\alpha-\beta)^{2}+4(d-\beta)$ has to be a perfect square. Then, it is straightforward to check that there are only three possible parameters $(n,d,\alpha,\beta)$ for $G$:
  • (a)
    $(n,\alpha+4,\alpha,\alpha)$;
  • (b)
    $(n,\alpha+3,\alpha,\alpha+1)$;
  • (c)
    $(n,\alpha+2,\alpha,\alpha+2)$.
Notice that $n$ can be uniquely determined by the other parameters $d,\alpha,\beta$. Indeed, we have by Theorem 2.6 that
$n-d-1=\frac{d(d-\alpha-1)}{\beta}.$ (4.3)
In the case of $(n,\alpha+4,\alpha,\alpha)$. Since $\frac{d(d-\alpha-1)}{\beta}=3+\frac{12}{\alpha}$ must be a positive integer, the possible values of $\alpha$ are $1$, $2$, $3$, $4$, $6$, $12$. By (4.3), we have $n=\alpha+\frac{12}{\alpha}+8$ and hence the multiplicity numbers $m_{2}=\frac{(4+\alpha)(6+\alpha)}{4\alpha}$ and $m_{3}=\frac{3(2+\alpha)(4+\alpha)}{4\alpha}$. Since $m_{2}$ and $m_{3}$ have to be positive integers, the possible values of $\alpha$ are $2$, $4$, $6$, $12$. Therefore, the possible parameters in this case are
  • (1)
    $(16,6,2,2)$;
  • (2)
    $(15,8,4,4)$;
  • (3)
    $(16,10,6,6)$;
  • (4)
    $(21,16,12,12)$.
In the case of $(n,\alpha+3,\alpha,\alpha+1)$. Since $\frac{d(d-\alpha-1)}{\beta}=2+\frac{4}{\alpha+1}$ must be a positive integer, the possible values of $\alpha$ are 00, $1$, $3$. Therefore, the possible parameters are
  • (5)
    $(10,3,0,1)$;
  • (6)
    $(9,4,1,2)$;
  • (7)
    $(10,6,3,4)$.
In the case of $(n,\alpha+2,\alpha,\alpha+2)$. By (4.3) and Theorem 4.1, we calculate that $m_{2}=\frac{4+\alpha}{2}$ and $m_{3}=\frac{2+\alpha}{2}$. Since $m_{2}$ and $m_{3}$ have to be positice integers, $\alpha$ must be even. Therefore, the possible parameters are
  • (8)
    $(2k,2k-2,2k-4,2k-2),k\geq 2$.
By Theorem 4.1, it is direct to check for all possible graphs $(1)-(8)$ that
$\lambda_{2}=\frac{2+\alpha}{d}.$
This completes the proof. ∎
Proof of Theorem 1.7.
Since $G$ is Lichnerowicz sharp, we have
$\lambda_{2}=\min_{xy\in E}\kappa(x,y)\leq\frac{2+\alpha}{d},$
where the last inequality is due to Theorem 1.1. By Theorem 4.2, we conclude that $\min_{xy\in E}\kappa(x,y)=\lambda_{2}=\frac{2+\alpha}{d}$. Therefore, we obtain $\kappa(x,y)=\frac{2+\alpha}{d}$ for all $xy\in E$ by applying Theorem 1.1 again. This completes the proof. ∎
We recall the following classificaiton result from [14].
Theorem 4.3.
[14, Theorems 6.3 and 6.4] The Lichnerowicz sharp strongly regular graphs with smallest adjacency eigenvalue $-2$ are precisely the following ones: The cocktail party graphs $CP(n)$, $n\geq 2$, the lattice graphs $L_{2}(n)\cong K_{n}\times K_{n}$, $n\geq 3$, the triangular graphs $T(n)\cong J(n,2)$, $n\geq 5$, the demi-cube $Q^{5}_{(2)}$, and the Schläfli graph.
Proof of Corollary 1.8.
By Theorem 4.1, any strongly regular graph with $\lambda_{2}=\frac{2+\alpha}{d}$ has the smallest adjacency eigenvalue $-2$. Thus, by Theorem 4.3, we determine all the Lichnerowicz sharp strongly regular graphs with $\beta\geq\alpha$ as following: The cocktail party graphs $CP(k),k\geq 2$, the Paley graph $P(9)$, the Johnson graphs $J(5,2)$ and $J(6,2)$ and the demi-cube $Q^{5}_{(2)}$. ∎
Acknowledgement
This work has been supported by the National Key R & D Program of China 2023YFA1010200 and the National Natural Science Foundation of China No. 12031017 and No. 12431004. H. Z. is very grateful to Professor Wen Huang for his encouragement and helpful discussion.
Appendix
In this appendix, we present two lists of parameters $(n,d,\alpha,\beta)$ for strongly regular graphs with at most $512$ vertices based on the data in [9, Chapter 12]: Table LABEL:table:1234 provides strongly regular graphs in the list of [9, Chapter 12] that fulfill at least one of the conditions (1), (2), (3), and (4), and particularly, $\beta>\alpha$; Table LABEL:table:5 includes strongly regular graphs in the list of [9, Chapter 12] that meet the condition (5), and particularly $\beta\leq\alpha$.
We first explain the notations column by column following [9, Chapter 12].
  • (i)
    Existence: A number indicates the precise number of non-isomorphic examples. "!" when there is a unique such graph, "+" when there is a known example, and "?" when there is not a known example.
  • (ii)
    The parameters $(n,d,\alpha,\beta)$: the number of vertices, the degree, the number of common neighbors of two adjacent vertices, and the number of common neighbors of two nonadjacent vertices, respectively.
  • (iii)
    The smallest non-zero Laplacian eigenvalue $\lambda_{2}$ and the Lin-Lu-Yau curvature $\kappa$.
  • (iv)
    Comments. The symbol ‘↓’ labels a descendant of a regular $2$-graph on $v+1$ vertices. The symbol ‘↑’ labels a graph in the switching class of a regular $2$-graph. Labels are postfixed with ‘?’ when the corresponding object is unknown.
Table 1: Some strongly regular graphs satisfying one of conditions (1)-(4)
ex n d $\alpha$ $\beta$ $\lambda_{2}$ $k(x,y)$ comment
15! 25 12 5 6 $\frac{5}{6}$ $\frac{7}{12}$ Paulus-Rozenfel’d; OA(5, 3); Paley(25); ↓
41! 29 14 6 7 $\frac{1181}{1400}$ $\frac{4}{7}$ Enumerated by Bussemaker and by Spence; Paley(29); ↓
3854! 35 16 6 8 $\frac{7}{8}$ $\frac{1}{2}$ Enumerated by McKay & Spence; no pg(5, 4, 2) (De Clerck); ↓
+ 37 18 8 9 $\frac{773}{900}$ $\frac{5}{9}$ see McKay & Spence; Crnković & Maksimović ; Maksimović & Rukavina; Paley(37); ↓
+ 41 20 9 10 $\frac{173}{200}$ $\frac{11}{20}$ Maksimović & Rukavina; Paley(41);
+ 45 22 10 11 $\frac{383}{440}$ $\frac{6}{11}$ Mathon; ↓
+ 49 24 11 12 $\frac{7}{8}$ $\frac{13}{24}$ Paley(49); OA(7, 4); ↓
+ 50 28 15 16 $\frac{25}{28}$ $\frac{17}{28}$ S(2, 4, 25); ↑
+ 53 26 12 13 $\frac{1143}{1300}$ $\frac{7}{13}$ Paley(53); ↓
+ 61 30 14 15 $\frac{2659}{3000}$ $\frac{8}{15}$ Paley(61); ↓
+ 63 30 13 15 $\frac{9}{10}$ $\frac{1}{2}$ qs $2-(36,16,12)$; $O_{7}(2)$; $Sp_{6}(2)$; pg(7, 5, 3); ↓
+ 64 35 18 20 $\frac{32}{35}$ $\frac{4}{7}$ OA(8, 5); Wallis; Goethals-Seidel; $VO^{+}_{6}(2)$; ↑
+ 65 32 15 16 $\frac{2847}{3200}$ $\frac{17}{32}$ Gritsenko; ↓
+ 73 36 17 18 $\frac{3223}{3600}$ $\frac{19}{36}$ Paley(73); ↓
+ 81 40 19 20 $\frac{9}{10}$ $\frac{21}{40}$ Paley(81); OA(9, 5); ↓
+ 82 45 24 25 $\frac{41}{45}$ $\frac{26}{45}$ S(2, 5, 41); ↑
? 85 42 20 21 $\frac{1263}{1400}$ $\frac{11}{21}$
+ 89 44 21 22 $\frac{1989}{2200}$ $\frac{23}{44}$ Paley(89); ↓
+ 97 48 23 24 $\frac{2179}{2400}$ $\frac{25}{48}$ Paley(97); ↓
+ 99 48 22 24 $\frac{11}{12}$ $\frac{1}{2}$ no pg(9, 6, 4)(Lam et al.); ↓
+ 100 54 28 30 $\frac{25}{27}$ $\frac{5}{9}$ OA(10, 6)?; ↑
+ 101 50 24 25 $\frac{1137}{1250}$ $\frac{13}{25}$ Paley(101); ↓
? 105 64 38 40 $\frac{15}{16}$ $\frac{5}{8}$
+ 109 54 26 27 $\frac{616}{675}$ $\frac{14}{27}$ Paley(109); ↓
+ 113 56 27 28 $\frac{2559}{2800}$ $\frac{29}{56}$ Paley(113); ↓
? 117 58 28 29 $\frac{5309}{5800}$ $\frac{15}{29}$?
? 120 84 58 60 $\frac{20}{21}$ $\frac{5}{7}$
+ 121 60 29 30 $\frac{11}{12}$ $\frac{31}{60}$ Paley(121); OA(11, 6); ↓
+ 122 66 35 36 $\frac{61}{66}$ $\frac{37}{66}$ S(2, 6, 61); ↑
+ 125 52 15 26 $\frac{25}{26}$ $\frac{17}{52}$ pg(5,13,2)?; ↓
+ 125 62 30 31 $\frac{5691}{6200}$ $\frac{16}{31}$ Paley(125); ↓
+ 137 68 33 34 $\frac{1253}{1360}$ $\frac{35}{68}$ Paley(137); ↓
+ 143 70 33 35 $\frac{13}{14}$ $\frac{1}{2}$ qs $2-(78,36,30)$; pg(11,7,5)?; ↓
+ 144 77 40 42 $\frac{72}{77}$ $\frac{6}{11}$ OA(12, 7); Wallis; Goethals-Seidel;
? 145 72 35 36 $\frac{277}{300}$ $\frac{37}{72}$?
+ 149 74 36 37 $\frac{171}{185}$ $\frac{19}{37}$ Paley(149); ↓
? 153 76 37 38 $\frac{879}{950}$ $\frac{39}{76}$?
? 154 72 26 40 $\frac{35}{36}$ $\frac{7}{18}$
+ 157 78 38 39 $\frac{301}{325}$ $\frac{20}{39}$ Paley(157); ↓
? 160 105 68 70 $\frac{20}{21}$ $\frac{2}{3}$
+ 169 84 41 42 $\frac{13}{14}$ $\frac{43}{84}$ Paley(169); OA(13,7); ↓
+ 170 91 48 49 $\frac{85}{91}$ $\frac{50}{91}$ S(2, 7, 85);
+ 173 86 42 43 $\frac{999}{1075}$ $\frac{22}{43}$ Paley(173); ↓
+ 175 72 20 36 $\frac{35}{36}$ $\frac{11}{36}$ Haemers pg(5,18,2); edges of Hoffman-Singleton; ↓
+ 176 90 38 54 $\frac{44}{45}$ $\frac{4}{9}$ pg(6,18,3)?; ↑
+ 181 90 44 45 $\frac{8377}{9000}$ $\frac{23}{45}$ Paley(181); ↓
? 185 92 45 46 $\frac{857}{920}$ $\frac{47}{92}$?
? 189 140 103 105 $\frac{27}{28}$ $\frac{3}{4}$
+ 193 96 47 48 $\frac{597}{640}$ $\frac{49}{96}$ Paley(193); ↓
+ 195 96 46 48 $\frac{15}{16}$ $\frac{1}{2}$ pg(13,8,6)?; ↓
+ 196 114 59 76 $\frac{56}{57}$ $\frac{61}{114}$ pg(7,19,4)?
+ 196 104 54 56 $\frac{49}{52}$ $\frac{7}{13}$ OA(14,8)?; ↑
+ 197 98 48 49 $\frac{2287}{2450}$ $\frac{25}{49}$ Paley(197); ↓
? 205 102 50 51 $\frac{1589}{1700}$ $\frac{26}{51}$?
? 210 132 82 84 $\frac{21}{22}$ $\frac{7}{11}$
? 221 110 54 55 $\frac{937}{1000}$ $\frac{28}{55}$?
+ 225 112 55 56 $\frac{15}{16}$ $\frac{57}{112}$ ConfMat$(16)^{2}$; OA(15, 8)?; ↓
+ 226 120 63 64 $\frac{113}{120}$ $\frac{13}{24}$ S(2, 8, 113);
+ 229 114 56 57 $\frac{10693}{11400}$ $\frac{29}{57}$ Paley(229); ↓
? 231 160 110 112 $\frac{77}{80}$ $\frac{7}{10}$
+ 233 116 57 58 $\frac{10887}{11600}$ $\frac{59}{116}$ Paley(233); ↓
+ 241 120 59 60 $\frac{1879}{2000}$ $\frac{61}{120}$ Paley(241); ↓
+ 243 110 37 60 $\frac{54}{55}$ $\frac{39}{110}$ Delsarte; $[55,5]_{3}$ (wts 36,45)
? 245 108 39 54 $\frac{35}{36}$ $\frac{41}{108}$ pg(7,18,3)?; ↓?
? 245 122 60 61 $\frac{11467}{12200}$ $\frac{31}{61}$?
+ 253 112 36 60 $\frac{55}{56}$ $\frac{19}{56}$ S(4, 7, 23); $M_{23}$
+ 255 126 61 63 $\frac{17}{18}$ $\frac{1}{2}$ $O_{9}(2)$; $Sp_{8}(2)$; pg(15, 9, 7); ↓
+ 256 135 70 72 $\frac{128}{135}$ $\frac{8}{15}$ OA(16, 9); Wallis; Goethals-Seidel; $VO^{+}_{8}(2)$; ↑
+ 257 128 63 64 $\frac{753}{800}$ $\frac{65}{128}$ Paley(257); ↓
? 261 130 64 65 $\frac{6121}{6500}$ $\frac{33}{65}$?
? 265 132 65 66 $\frac{3109}{3300}$ $\frac{67}{132}$?
+ 269 134 66 67 $\frac{1263}{1340}$ $\frac{34}{67}$ Paley(269); ↓
! 275 112 30 56 $\frac{55}{56}$ $\frac{2}{7}$ $q_{22}^{2}=0$; no pg(5, 28, 2)
+ 276 140 58 84 $\frac{69}{70}$ $\frac{3}{7}$ pg(6,28,3)?; ↑
+ 277 138 68 69 $\frac{283}{300}$ $\frac{35}{69}$ Paley(277); ↓
? 280 216 166 168 $\frac{35}{36}$ $\frac{7}{9}$
+ 281 140 69 70 $\frac{3303}{3500}$ $\frac{71}{140}$ Paley(281); ↓
? 287 126 45 63 $\frac{41}{42}$ $\frac{47}{126}$ pg(7,21,3)?; ↓?
? 288 182 106 130 $\frac{90}{91}$ $\frac{54}{91}$ pg(8,26,5)?
? 288 147 66 84 $\frac{48}{49}$ $\frac{68}{147}$ pg(8,21,4)?; ↑?
+ 289 144 71 72 $\frac{17}{18}$ $\frac{73}{144}$ Paley(289); OA(17,9); ↓
+ 290 153 80 81 $\frac{145}{153}$ $\frac{82}{153}$ S(2,9,145)?; ↑
+ 293 146 72 73 $\frac{6897}{7300}$ $\frac{37}{73}$ Paley(293); ↓
? 300 182 100 126 $\frac{90}{91}$ $\frac{51}{91}$
? 301 150 65 84 $\frac{49}{50}$ $\frac{67}{150}$
? 305 152 75 76 $\frac{14377}{15200}$ $\frac{77}{152}$?
+ 313 156 77 78 $\frac{2953}{3120}$ $\frac{79}{156}$ Paley(313); ↓
+ 317 158 78 79 $\frac{374}{395}$ $\frac{40}{79}$ Paley(317); ↓
? 320 231 166 168 $\frac{32}{33}$ $\frac{8}{11}$
+ 323 160 78 80 $\frac{19}{20}$ $\frac{1}{2}$ pg(17,10,8)?; ↓
+ 324 170 88 90 $\frac{81}{85}$ $\frac{9}{17}$ OA(18,10)?; ↑
? 325 252 195 196 $\frac{35}{36}$ $\frac{197}{252}$
? 325 162 80 81 $\frac{15349}{16200}$ $\frac{41}{81}$?
? 333 166 82 83 $\frac{7869}{8300}$ $\frac{42}{83}$?
? 336 200 118 120 $\frac{24}{25}$ $\frac{3}{5}$
+ 337 168 83 84 $\frac{569}{600}$ $\frac{85}{168}$ Paley(337); ↓
+ 343 150 53 75 $\frac{49}{50}$ $\frac{11}{30}$ Godsil; pg(7,25,3)?; ↓
+ 344 175 78 100 $\frac{172}{175}$ $\frac{16}{35}$ pg(8,25,4)?; Taylor ↑
+ 349 174 86 87 $\frac{4129}{4350}$ $\frac{44}{87}$ Paley(349); ↓
? 351 280 223 224 $\frac{39}{40}$ $\frac{45}{56}$
+ 351 224 142 144 $\frac{27}{28}$ $\frac{9}{14}$ $NO_{7}^{-}(3)$
? 351 210 113 144 $\frac{104}{105}$ $\frac{23}{42}$
+ 353 176 87 88 $\frac{16711}{17600}$ $\frac{89}{176}$ Paley(353); ↓
+ 361 180 89 90 $\frac{19}{20}$ $\frac{91}{180}$ Paley(361); OA(19,10); ↓
+ 362 190 99 100 $\frac{181}{190}$ $\frac{101}{190}$ S(2,10,181)?; ↑
? 365 182 90 91 $\frac{3459}{3640}$ $\frac{46}{91}$?
? 369 184 91 92 $\frac{1749}{1840}$ $\frac{93}{184}$?
+ 373 186 92 93 $\frac{4421}{4650}$ $\frac{47}{93}$ Paley(373); ↓
? 375 198 93 117 $\frac{65}{66}$ $\frac{95}{198}$
? 377 188 93 94 $\frac{17879}{18800}$ $\frac{95}{188}$?
? 378 260 178 180 $\frac{63}{65}$ $\frac{9}{13}$ $NO_{7}^{+}(3)$
+ 389 194 96 97 $\frac{2308}{2425}$ $\frac{49}{97}$ Paley(389); ↓
+ 397 198 98 99 $\frac{857}{900}$ $\frac{50}{99}$ Paley(397); ↓
+ 399 198 97 99 $\frac{21}{22}$ $\frac{1}{2}$ pg(19,11,9)?; ↓
+ 400 209 108 110 $\frac{200}{209}$ $\frac{10}{19}$ OA(20,11)?; Wallis; Goethals-Seidel; ↑
+ 401 200 99 100 $\frac{19049}{20000}$ $\frac{101}{200}$ Paley(401); ↓
? 405 202 100 101 $\frac{4811}{5050}$ $\frac{51}{101}$?
+ 409 204 101 102 $\frac{19439}{20400}$ $\frac{103}{204}$ Paley(409); ↓
+ 421 210 104 105 $\frac{2503}{2625}$ $\frac{53}{105}$ Paley(421); ↓
? 425 264 163 165 $\frac{85}{88}$ $\frac{5}{8}$
? 425 212 105 106 $\frac{20219}{21200}$ $\frac{107}{212}$?
? 429 320 238 240 $\frac{39}{40}$ $\frac{3}{4}$
+ 433 216 107 108 $\frac{229}{240}$ $\frac{109}{216}$ Paley(433); ↓
? 441 220 95 124 $\frac{217}{220}$ $\frac{97}{220}$
+ 441 220 109 110 $\frac{21}{22}$ $\frac{111}{220}$ Mathon; OA(21,11)?; ↓
+ 442 231 120 121 $\frac{221}{231}$ $\frac{122}{231}$ S(2,11,221)?; ↑?
? 445 222 110 111 $\frac{1413}{1480}$ $\frac{56}{111}$?
? 448 297 196 198 $\frac{32}{33}$ $\frac{2}{3}$
+ 449 224 111 112 $\frac{21391}{22400}$ $\frac{113}{224}$ Paley(449); ↓
+ 457 228 113 114 $\frac{21781}{22800}$ $\frac{115}{228}$ Paley(457); ↓
? 459 208 82 104 $\frac{51}{52}$ $\frac{21}{52}$ pg(9,26,4)?; ↓?
? 460 243 114 144 $\frac{80}{81}$ $\frac{116}{243}$
+ 461 230 114 115 $\frac{2747}{2875}$ $\frac{58}{115}$ Paley(461); ↓
? 477 238 118 119 $\frac{11379}{11900}$ $\frac{60}{119}$?
? 481 240 119 120 $\frac{7651}{8000}$ $\frac{121}{240}$?
? 483 240 118 120 $\frac{23}{24}$ $\frac{1}{2}$ pg(21,12,10)?; ↓?
? 484 252 130 132 $\frac{121}{126}$ $\frac{11}{21}$ OA(22,12)?; ↑?
? 485 242 120 121 $\frac{23149}{24200}$ $\frac{61}{121}$?
? 493 246 122 123 $\frac{1177}{1230}$ $\frac{62}{123}$?
? 495 234 93 126 $\frac{77}{78}$ $\frac{95}{234}$
? 505 252 125 126 $\frac{4021}{4200}$ $\frac{127}{252}$?
? 507 276 135 168 $\frac{91}{92}$ $\frac{137}{276}$
+ 509 254 126 127 $\frac{12161}{12700}$ $\frac{64}{127}$ Paley(509); ↓
Table 2: Some strongly regular graphs satisfying the condition (5)
ex n d $\alpha$ $\beta$ $\lambda_{2}$ $k(x,y)$ comment
28! 40 27 18 18 $\frac{8}{9}$ $\frac{20}{27}$ $NU_{4}(2)$
! 56 45 36 36 $\frac{14}{15}$ $\frac{38}{45}$ qs 2-(21,6,4)
167! 64 45 32 30 $\frac{8}{9}$ $\frac{34}{45}$
! 77 60 47 45 $\frac{11}{12}$ $\frac{49}{60}$ Witt: qs 2-(22,6,5)
! 81 60 45 42 $\frac{9}{10}$ $\frac{47}{60}$
+ 85 64 48 48 $\frac{15}{16}$ $\frac{25}{32}$
? 88 60 41 40 $\frac{11}{12}$ $\frac{43}{60}$
+ 96 76 60 60 $\frac{18}{19}$ $\frac{31}{38}$
! 100 77 60 56 $\frac{10}{11}$ $\frac{62}{77}$ $q_{11}^{1}=0$
? 111 80 58 56 $\frac{37}{40}$ $\frac{3}{4}$
! 112 81 60 54 $\frac{8}{9}$ $\frac{62}{81}$ $q_{11}^{1}=0$
? 112 75 50 50 $\frac{14}{15}$ $\frac{52}{75}$
? 115 96 80 80 $\frac{23}{24}$ $\frac{41}{48}$
? 120 85 60 60 $\frac{16}{17}$ $\frac{62}{85}$
? 121 84 59 56 $\frac{11}{12}$ $\frac{61}{84}$
+ 125 96 74 72 $\frac{15}{16}$ $\frac{19}{24}$
? 133 100 75 75 $\frac{19}{20}$ $\frac{77}{100}$
+ 144 104 76 72 $\frac{12}{13}$ $\frac{3}{4}$
? 154 105 72 70 $\frac{14}{15}$ $\frac{74}{105}$
+ 156 125 100 100 $\frac{24}{25}$ $\frac{102}{125}$
? 162 140 121 120 $\frac{27}{28}$ $\frac{123}{140}$
+ 162 128 100 96 $\frac{15}{16}$ $\frac{51}{64}$
? 169 126 95 90 $\frac{13}{14}$ $\frac{97}{126}$
? 171 120 84 84 $\frac{19}{20}$ $\frac{43}{60}$
? 176 150 128 126 $\frac{24}{25}$ $\frac{13}{15}$
+ 176 126 90 90 $\frac{20}{21}$ $\frac{46}{63}$
? 183 130 93 90 $\frac{61}{65}$ $\frac{19}{26}$
? 196 156 125 120 $\frac{49}{52}$ $\frac{127}{156}$
? 196 150 116 110 $\frac{14}{15}$ $\frac{59}{75}$
+ 196 135 94 90 $\frac{14}{15}$ $\frac{32}{45}$
? 204 175 150 150 $\frac{34}{35}$ $\frac{152}{175}$
? 208 162 126 126 $\frac{26}{27}$ $\frac{64}{81}$
? 210 176 148 144 $\frac{21}{22}$ $\frac{75}{88}$
+ 216 175 142 140 $\frac{24}{25}$ $\frac{144}{175}$
? 217 150 105 100 $\frac{14}{15}$ $\frac{107}{150}$
? 220 147 98 98 $\frac{20}{21}$ $\frac{100}{147}$
? 225 176 139 132 $\frac{15}{16}$ $\frac{141}{176}$
? 225 160 115 110 $\frac{15}{16}$ $\frac{117}{160}$
? 232 198 169 168 $\frac{32}{33}$ $\frac{19}{22}$
? 232 168 122 120 $\frac{20}{21}$ $\frac{31}{42}$
? 235 182 141 140 $\frac{25}{26}$ $\frac{11}{14}$
? 238 162 111 108 $\frac{17}{18}$ $\frac{113}{162}$
? 243 176 130 120 $\frac{81}{88}$ $\frac{3}{4}$
? 245 192 152 144 $\frac{15}{16}$ $\frac{77}{96}$
? 250 168 113 112 $\frac{20}{21}$ $\frac{115}{168}$
+ 256 204 164 156 $\frac{16}{17}$ $\frac{83}{102}$ vanLint-Schrijver
+ 256 187 138 132 $\frac{16}{17}$ $\frac{140}{187}$
? 259 216 180 180 $\frac{35}{36}$ $\frac{91}{108}$
? 260 189 138 135 $\frac{20}{21}$ $\frac{20}{27}$
? 261 196 147 147 $\frac{27}{28}$ $\frac{149}{196}$
? 266 220 183 176 $\frac{19}{20}$ $\frac{37}{44}$
? 273 192 136 132 $\frac{91}{96}$ $\frac{23}{32}$
? 276 200 148 136 $\frac{23}{25}$ $\frac{3}{4}$
? 276 200 145 144 $\frac{24}{25}$ $\frac{147}{200}$
? 280 217 168 168 $\frac{30}{31}$ $\frac{170}{217}$
? 285 220 171 165 $\frac{19}{20}$ $\frac{173}{220}$
? 288 246 210 210 $\frac{40}{41}$ $\frac{106}{123}$
? 289 216 163 156 $\frac{17}{18}$ $\frac{55}{72}$
? 289 198 137 132 $\frac{17}{18}$ $\frac{139}{198}$
+ 300 234 183 180 $\frac{25}{26}$ $\frac{185}{234}$
? 306 250 205 200 $\frac{24}{25}$ $\frac{207}{250}$
? 306 245 196 196 $\frac{34}{35}$ $\frac{198}{245}$
? 320 232 168 168 $\frac{28}{29}$ $\frac{85}{116}$
? 322 225 160 150 $\frac{14}{15}$ $\frac{18}{25}$
? 324 255 202 195 $\frac{81}{85}$ $\frac{4}{5}$
? 324 247 190 182 $\frac{18}{19}$ $\frac{192}{247}$
? 324 228 162 156 $\frac{18}{19}$ $\frac{41}{57}$
? 325 270 225 220 $\frac{26}{27}$ $\frac{227}{270}$
+ 325 256 204 192 $\frac{15}{16}$ $\frac{103}{128}$ $q_{11}^{1}=0$
? 329 288 252 252 $\frac{47}{48}$ $\frac{127}{144}$
? 340 231 158 154 $\frac{20}{21}$ $\frac{160}{231}$
? 341 256 192 192 $\frac{31}{32}$ $\frac{97}{128}$
+ 343 288 242 240 $\frac{35}{36}$ $\frac{61}{72}$
? 343 240 171 160 $\frac{14}{15}$ $\frac{173}{240}$
? 352 325 300 300 $\frac{64}{65}$ $\frac{302}{325}$
? 352 315 282 280 $\frac{44}{45}$ $\frac{284}{315}$
? 361 280 219 210 $\frac{19}{20}$ $\frac{221}{280}$
? 361 260 189 182 $\frac{19}{20}$ $\frac{191}{260}$
? 364 275 210 200 $\frac{52}{55}$ $\frac{212}{275}$
+ 364 243 162 162 $\frac{26}{27}$ $\frac{164}{243}$
? 375 308 253 252 $\frac{75}{77}$ $\frac{255}{308}$
? 375 264 188 180 $\frac{125}{132}$ $\frac{95}{132}$
+ 378 325 280 275 $\frac{63}{65}$ $\frac{282}{325}$
+ 378 261 180 180 $\frac{28}{29}$ $\frac{182}{261}$
? 381 266 187 182 $\frac{127}{133}$ $\frac{27}{38}$
? 385 324 273 270 $\frac{35}{36}$ $\frac{275}{324}$
? 392 345 304 300 $\frac{112}{115}$ $\frac{102}{115}$
+ 400 343 294 294 $\frac{48}{49}$ $\frac{296}{343}$
? 400 315 250 240 $\frac{20}{21}$ $\frac{4}{5}$
? 400 294 218 210 $\frac{20}{21}$ $\frac{110}{147}$
? 400 273 188 182 $\frac{20}{21}$ $\frac{190}{273}$
? 405 308 235 231 $\frac{27}{28}$ $\frac{237}{308}$
? 408 297 216 216 $\frac{32}{33}$ $\frac{218}{297}$
? 428 315 234 225 $\frac{20}{21}$ $\frac{236}{315}$
? 430 294 203 196 $\frac{20}{21}$ $\frac{205}{294}$
? 437 336 260 252 $\frac{23}{24}$ $\frac{131}{168}$
? 441 352 283 272 $\frac{21}{22}$ $\frac{285}{352}$
? 441 330 249 240 $\frac{21}{22}$ $\frac{251}{330}$
? 441 308 217 210 $\frac{21}{22}$ $\frac{219}{308}$
? 451 320 228 224 $\frac{77}{80}$ $\frac{23}{32}$
? 456 375 310 300 $\frac{24}{25}$ $\frac{104}{125}$
? 456 351 270 270 $\frac{38}{39}$ $\frac{272}{351}$
? 456 325 236 220 $\frac{304}{325}$ $\frac{238}{325}$
? 456 315 218 216 $\frac{304}{315}$ $\frac{44}{63}$
? 460 360 282 280 $\frac{35}{36}$ $\frac{71}{90}$
? 460 312 213 208 $\frac{23}{24}$ $\frac{215}{312}$
? 465 320 220 220 $\frac{31}{32}$ $\frac{111}{160}$
? 470 343 252 245 $\frac{47}{49}$ $\frac{254}{343}$
? 477 336 240 228 $\frac{53}{56}$ $\frac{121}{168}$
? 484 391 318 306 $\frac{22}{23}$ $\frac{320}{391}$
? 484 378 297 288 $\frac{121}{126}$ $\frac{299}{378}$
? 484 368 282 272 $\frac{22}{23}$ $\frac{71}{92}$
? 484 345 248 240 $\frac{22}{23}$ $\frac{50}{69}$
? 486 388 310 308 $\frac{189}{194}$ $\frac{78}{97}$
? 490 345 248 230 $\frac{14}{15}$ $\frac{50}{69}$
? 494 408 337 336 $\frac{133}{136}$ $\frac{113}{136}$
? 495 456 420 420 $\frac{75}{76}$ $\frac{211}{228}$
? 496 441 392 392 $\frac{62}{63}$ $\frac{394}{441}$
? 496 385 300 294 $\frac{372}{385}$ $\frac{302}{385}$
? 505 420 351 340 $\frac{101}{105}$ $\frac{353}{420}$
? 506 405 324 324 $\frac{44}{45}$ $\frac{326}{405}$
? 507 462 421 420 $\frac{65}{66}$ $\frac{141}{154}$
? 507 352 246 240 $\frac{169}{176}$ $\frac{31}{44}$
? 511 432 366 360 $\frac{35}{36}$ $\frac{23}{27}$
+ 512 441 380 378 $\frac{48}{49}$ $\frac{382}{441}$
+ 512 378 282 270 $\frac{20}{21}$ $\frac{142}{189}$

### v3

@@
 Abstract
-We confirm a conjecture of Bonini et al.
-on the precise Lin–Lu–Yau curvature values of conference graphs, i.e., strongly regular graphs with parameters $(4\gamma+1,2\gamma,\gamma-1,\gamma)$, with $\gamma\geq 2$.
-Our method only depends on the parameter relations and applies to more general classes of amply regular graphs.
-In particular, we develop a new combinatorial method for showing the existence of local perfect matchings.
+We prove a conjecture of Bonini et al.
+on the precise values of the Lin–Lu–Yau curvature of conference graphs, i.e., strongly regular graphs with parameters $(4\gamma+1,2\gamma,\gamma-1,\gamma)$.
+Our method depends only on the parameter relations and applies to broader classes of amply regular graphs.
+In particular, we develop a new combinatorial approach to show the existence of local perfect matchings.
 A key observation is that counting common neighbors leads to useful quadratic polynomials.
-Our result also leads to an interesting number theoretic consequence on quadratic residues.
+As a corollary, we derive an interesting number-theoretic result concerning quadratic residues.
 1 Introduction
@@
 In [21], Lin, Lu, and Yau introduced a modified version of Ollivier’s coarse Ricci curvature on graphs.
-This concept of modified curvature, which we call the Lin–Lu–Yau curvature, has been proved to be very useful to detect geometric and analytic properties of the underlying graphs [14, 13, 12, 25, 26, 27].
-Explicit calculations of Lin–Lu–Yau curvature on graphs with local regularity restrictions have important applications in the study of locally correctable and some locally testable binary linear codes [20], and in bounding the diameter and eigenvalues of amply regular graphs [22, 19].
-A $d$-regular graph $G$ with $n$ vertices is called an amply regular graph with parameters $(n,d,\alpha,\beta)$ if any two adjacent vertices have $\alpha$ common neighbors and any two vertices of distance $2$ have $\beta$ common neighbors.
-Amply regular graphs with diameter $2$ are strongly regular.
-For any two adjacent vertices, denote by $\kappa(x,y)$ the Lin–Lu–Yau curvature of the edge $xy$.
-In the works of [22, 19], it has been shown that for any two adjacent vertices $x,y$ of an amply regular graph $G$ with parameters $(n,d,\alpha,\beta)$, the following lower bound estimates hold:
-  • If $\beta=\alpha>1$, then $\kappa(x,y)\geq\frac{2}{d}$.
-  • If $\beta>\alpha\geq 1$, then $\kappa(x,y)\geq\frac{3}{d}$.
-The Lin–Lu–Yau curvature of amply regular graphs with $\alpha>\beta\geq 2$ is rather mysterious. It is still open whether such graphs have positively lower bounded Lin–Lu–Yau curvature or not.
+This modified curvature, which we call the Lin–Lu–Yau curvature, has proved very useful in detecting geometric and analytic properties of the underlying graphs [12, 13, 14, 23, 24, 25, 28].
+Explicit calculations of Lin–Lu–Yau curvature on graphs with local regularity restrictions have important applications in the study of locally correctable and some locally testable binary linear codes [18], and in bounding the diameter and eigenvalues of amply regular graphs [17].
+A $d$-regular graph $G$ with $n$ vertices is called an amply regular graph with parameters $(n,d,\alpha,\beta)$ if any two adjacent vertices have $\alpha$ common neighbors and any two vertices at distance $2$ have $\beta$ common neighbors.
+Amply regular graphs with diameter $2$ are called strongly regular graphs.
+Denote by $\kappa(x,y)$ the Lin–Lu–Yau curvature of an edge $xy$.
+It is shown [17] that, for any edge $xy$ of an amply regular graph $G$ with parameters $(n,d,\alpha,\beta)$ with $\beta>\alpha\geq 1$, we have $\kappa(x,y)\geq\frac{3}{d}$.
 On the other hand, the Lin–Lu–Yau curvature has a natural upper bound.
@@
 Theorem 1.1.
-Let $G=(V,E)$ be a $d$-regular graph. For any edge $xy\in E$, we have
+[14] Let $G=(V,E)$ be a $d$-regular graph. For any edge $xy\in E$, we have
 $\kappa(x,y)\leq\frac{2+|\Delta_{xy}|}{d},$
@@
     there is a perfect matching between $N_{x}$ and $N_{y}$, where $N_{x}:=\Gamma(x)\backslash(\Delta_{xy}\cup\{y\})$ and $N_{y}:=\Gamma(y)\backslash(\Delta_{xy}\cup\{x\})$.
-The crucial role of local matching conditions in the curvature calculation has been first observed in [3, 32], and further developed in [5, 15, 27].
-For an amply regular graph $G$ with parameters $(n,d,\alpha,\beta)$, the above upper bound reads $\kappa(x,y)\leq\frac{2+\alpha}{d}$.
+The crucial role of local matching conditions in the curvature calculation was first observed in [3, 29], and further developed in [5, 15, 25].
+For an amply regular graph with parameters $(n,d,\alpha,\beta)$, the above upper bound reads $\kappa(x,y)\leq\frac{2+\alpha}{d}$.
 A natural question is for which parameters this upper bound is achieved.
 Bonini et al.
-[5] derived formulas for the Lin–Lu–Yau curvature of strongly regular graphs based on the parameters of graphs and the size of a maximal matching in the core neighborhood.
+[5] derived formulas for the Lin–Lu–Yau curvature of strongly regular graphs in terms of the graph parameters and the size of a maximal matching in the core neighborhood.
 They further proposed the following conjecture [5, Conjecture 1.7].
 Conjecture 1.2.
-Let $G=(V,E)$ be a strongly regular conference graph with parameters $(4\gamma+1,2\gamma,\gamma-1,\gamma)$, with $\gamma\geq 2$. Then the Lin–Lu–Yau curvature satisfies
+[5] Let $G=(V,E)$ be a strongly regular conference graph with parameters $(4\gamma+1,2\gamma,\gamma-1,\gamma)$ with $\gamma\geq 2$. Then the Lin–Lu–Yau curvature satisfies
 $\kappa(x,y)=\frac{1}{2}+\frac{1}{2\gamma},\,\,\text{for all}\,\,xy\in E.$
 Recall that a conference graph is a strongly regular graph with parameters $(4\gamma+1,2\gamma,\gamma-1,\gamma)$.
-Conference graphs have very interesting properties: While being highly symmetric, they are quasirandom [1, Theorem 9.3.1].
-The existence of conference graphs are known for small values of $\gamma$ and for each $\gamma$ such that $14\gamma+$ 1 is a prime power (Paley graphs).
+Conference graphs have very interesting properties: While being highly symmetric, they are quasirandom [1].
+The existence of conference graphs is known for each $\gamma$ such that $4\gamma+1$ is a prime power (i.e.
+Paley graphs).
 Let $n=4\gamma+1$ be a prime power and $GF(n)$ be the unique finite field of order $n$.
-Recall that a Paley graph $P(n)$ of order $n$ is a graph on $n$ vertices with two vertices adjacent if their difference is a non-zero square in $GF(n)$.
-However, there are still many values of $\gamma$ for which the existence of conference graphs is open.
-Observe by Theorem 1.1 that the Lin–Lu–Yau curvature of a conference graph is no greater than $\frac{2+(\gamma-1)}{2\gamma}=\frac{1}{2}+\frac{1}{2\gamma}$.
-Therefore, Conjecture 1.2 claims the upper bound is achieved for conference graphs.
-We stress that, in general, the Lin–Lu–Yau curvature of an amply regular graph is not necessarily uniquely determined by its parameters $(n,d,\alpha,\beta)$.
-For instance, the $4\times 4$ Rook’s graph and the Shrikhande graph are both strongly regular with parameters $(16,6,2,2)$, but Bonini et al.
-[5] calculated their Lin–Lu–Yau curvatures to be $\kappa=\frac{2}{3}$ and $\kappa=\frac{1}{3}$, respectively.
-In particular, the $4\times 4$ Rook’s graph takes the upper bound of the Lin–Lu–Yau curvature, while the Shrikhande graph does not.
-Huang, Liu, and Xia showed [19, Proposition 3.6] that $\kappa(x,y)=\frac{2+\alpha}{d}$ for any amply regular graph with parameters $(n,d,\alpha,\beta)$ such that $d\leq 2\beta-\alpha-1$.
-However, the parameters of conference graphs violate the restriction $d\leq 2\beta-\alpha-1$.
-The lower curvature bound of amply regular graph with $\beta>\alpha\geq 1$ in [19], which has been mentioned earlier, applies to conference graphs and provides a positive lower bound $\frac{3}{d}$.
-After finishing the first version of this paper, we notice from a very recent preprint [6] that the Lin–Lu–Yau curvature of the special case of Paley graphs $P(n)$ of even power order $n=p^{2m}$ has been confirmed to be $\frac{2+\alpha}{d}$ in an unpublished manuscript by Bonini, Liu, Semmens and Tran.
-Very interesting extensions to generalized Paley graphs have been carried out in [6].
-Our main result in this paper confirms the Conjecture 1.2.
+Recall that a Paley graph $P(n)$ of order $n$ is a graph on $n$ vertices with two vertices adjacent iff their difference is a non-zero square in $GF(n)$.
+By Theorem 1.1, the Lin–Lu–Yau curvature of a conference graph is no greater than $\frac{2+(\gamma-1)}{2\gamma}=\frac{1}{2}+\frac{1}{2\gamma}$.
+Therefore, Conjecture 1.2 in fact claims that the upper bound is achieved for conference graphs.
+We stress that, in general, the Lin–Lu–Yau curvature of an amply regular graph is not necessarily uniquely determined by its parameters.
+For instance, the $4\times 4$ Rook’s graph and the Shrikhande graph are both strongly regular with parameters $(16,6,2,2)$.
+However, Bonini et al.
+[5] computed their Lin–Lu–Yau curvatures to be $\frac{2}{3}$ and $\frac{1}{3}$, respectively.
+In particular, the $4\times 4$ Rook’s graph attains the upper bound of the Lin–Lu–Yau curvature, while the Shrikhande graph does not.
+Huang, Liu, and Xia showed [17, Proposition 3.6] that $\kappa(x,y)=\frac{2+\alpha}{d}$ for any amply regular graph with parameters $(n,d,\alpha,\beta)$ such that $d\leq 2\beta-\alpha-1$.
+However, the parameters of conference graphs, namely $(4\gamma+1,2\gamma,\gamma-1,\gamma)$, do not satisfy this condition.
+A recent work by Bonini, Liu, Semmens, and Tran [6] proves Conjecture 1.2 for Paley graphs $P(n)$ of even power order $n=p^{2m}$.
+Very interesting extensions to generalized Paley graphs have also been carried out in [6].
+In this paper, we provide several sufficient conditions for the curvature of an amply regular graph to reach the natural upper bound. As a corollary, we prove Conjecture 1.2.
+Theorem 1.3.
+Let $G=(V,E)$ be an amply regular graph with parameters $(n,d,\alpha,\beta)$ such that $n<3d-2\alpha$. Suppose that one of the following conditions holds:
+  • (1)
+    $\beta=\alpha+1$ and $d>\alpha+\sqrt{6\alpha+\frac{1}{4}}+\frac{3}{2}$;
+  • (2)
+    $\beta\geq\alpha+2$ and $d\geq\frac{3}{2}\beta+\frac{1}{2}\sqrt{4\alpha^{2}-3\beta^{2}+28\beta}$;
+  • (3)
+    $\beta\geq\frac{2\sqrt{3}}{3}\alpha+7$.
+Then the Lin–Lu–Yau curvature satisfies
+$\kappa(x,y)=\frac{2+\alpha}{d},\,\,\text{for all}\,\,xy\in E.$
+In particular, condition (1) of Theorem 1.3 covers all conference graphs with $\gamma>6$. The cases for $2\leq\gamma\leq 6$ are verified separately in Section 3.
 Theorem 1.4.
 The Conjecture 1.2 holds true.
-Indeed, our method depends only on the parameter relations and applies to more general classes of amply regular graphs. Below, we provide several sufficient parameter conditions.
-Theorem 1.4.
-Let $G=(V,E)$ be an amply regular graph with parameters $(n,d,\alpha,\beta)$. Suppose that one of the following conditions holds:
-  • (1)
-    $\beta=\alpha+1$, $n<3d-2\alpha$ and $d>\alpha+\sqrt{6\alpha+\frac{1}{4}}+\frac{3}{2}$;
-  • (2)
-    $\beta=\alpha+2$, $n<3d-2\alpha$ and $d\geq\alpha+\sqrt{10\alpha+\frac{41}{4}}+\frac{3}{2}$;
-  • (3)
-    $\beta>\alpha+2$, $n<3d-2\alpha$ and $\left|d-\frac{3}{2}\beta\right|\geq\frac{1}{2}\sqrt{4\alpha^{2}-3\beta^{2}+4\alpha+24\beta-20}$;
-  • (4)
-    $\beta\geq\frac{2\sqrt{3}}{3}\alpha+7$ and $n<3d-2\alpha$;
-  • (5)
-[…]
-Then the Lin–Lu–Yau curvature satisfies
-$\kappa(x,y)=\frac{2+\alpha}{d},\,\,\text{for all}\,\,xy\in E.$
-By Theorem 1.1, Theorem 1.4 is equivalent to the following result in terms of local matchings.
+By Theorem 1.1, proving Theorem 1.3 reduces to establishing the following result in terms of local matchings, which is equivalent to Theorem 1.3.
 Theorem 1.3.
 Let $G=(V,E)$ be an amply regular graph with parameters $(n,d,\alpha,\beta)$.
-Suppose that one of the $5$ parameter conditions in Theorem 1.4 holds.
-Then there exists a perfect matching between $N_{x}$ and $N_{y}$ for all $xy\in E$.
-Interestingly, the parameter conditions stated in Theorem 1.4 are fulfilled by many known strongly regular graphs, such as Paley graphs, McLaughlin graph and so on.
-We provide a list of more graphs (on at most $512$ vertices) satisfying one of the parameter conditions in Theorem 1.4 in the Appendix.
-In particular, by the condition $(1)$ of Theorem 1.4, Conjecture 1.2 holds true for $\gamma>6$.
-The condition $(5)$ applies to many graphs of the mysterious case of amply regular graphs with $\alpha>\beta$, see the Appendix.
-Our proof provides a new combinatorial method for showing the existence of local perfect matchings.
-This new trick can also be employed to deal with many cases of amply regular graphs other than those in Theorem 1.4.
+If one of the $3$ parameter conditions in Theorem 1.3 holds, then there exists a perfect matching between $N_{x}$ and $N_{y}$ for every $xy\in E$.
+Besides conference graphs, our parameter conditions are also fulfilled by many other strongly regular graphs, as shown in the arXiv version of this paper (see [9, Appendix]).
+In fact, we provide five more refined conditions in [9, Theorem 1.4], which generalize the three conditions in Theorem 1.3.
+For brevity, we do not include them here.
+Our proof provides a new combinatorial method for showing the existence of local perfect matchings, which is also applicable to many cases of amply regular graphs beyond those in Theorem 1.3 and [9, Theorem 1.4].
 For example, the restriction $n<3d-2\alpha$ is not always necessary.
 We refer to Remark 3.4 and Example 3.5 for more detailed discussions.
-Applying Theorem 1.3 to Paley graphs, we obtain an interesting result of number theory.
+Applying Theorem 1.4 to Paley graphs yields an interesting number-theoretic result.
 Theorem 1.4 tells that there exists a perfect matching between $N_{x}$ and $N_{y}$ for any adjacent vertices $x,y$ of a Paley graph $P(n)$ with $n>5$.
-This leads to the following consequence.
+This leads to the following corollary.
 Corollary 1.5.
 Let $n>5$ be a prime power congruent to $1$ modulo $4$.
-Consider two numbers $x,y\in GF(n)$ such that $x-y$ is a non-zero square in $GF(n)$.
-For any subset $S\subset GF(n)\setminus\{x,y\}$ with $|S|\geq\frac{3}{4}(n-1)$, there exist $w,z\in S$ such that $x-w,w-z,z-y$ are non-zero squares and $x-z,y-w$ are not squares in $GF(n)$.
-In particular, this corollary applies to quadratic residues when $n>5$ is a Pythagorean prime, i.e., a prime congruent to $1$ modulo $4$.
-Furthermore, we study another upper bound of the Lin–Lu–Yau curvature, which is given in terms of the Laplacaian eigenvalue and is an analogue to the Lichnerowicz Theorem in Riemannian geometry.
-Theorem 1.6 (Discrete Lichnerowicz Theorem [28, 23]).
-Let $G=(V,E)$ be a finite connected graph. Then the Lin–Lu–Yau curvature statisfies
-$\min_{xy\in E}\kappa(x,y)\leq\lambda_{2},$
-where $\lambda_{2}$ is the first nonzero eigenvalue of the normalized Laplacian $L$.
-Here the normalized Laplacian $L$ is defined as $L=I-D^{-\frac{1}{2}}AD^{-\frac{1}{2}}$, with $A$ being the adjacency matrix of the graph $G$ and $D$ being the diagonal matrix of vertex degrees.
-Following [14], we say a graph $G$ is Lichnerowicz sharp if $\lambda_{2}=\min_{xy\in E}\kappa(x,y)$, i.e., the above upper bound is achieved. We observe the following interesting phenomena.
-Theorem 1.7.
-Let $G=(V,E)$ be a connected strongly regular graph with parameters $(n,d,\alpha,\beta)$ such that $\beta\geq\alpha$. If $G$ is Lichnerowicz sharp, i.e., $\min_{xy\in E}\kappa(x,y)=\lambda_{2}$, then
-$\kappa(x,y)=\frac{2+\alpha}{d}.$
-Indeed, all distance regular graphs with the two curvature upper bounds in Theorem 1.1 and Theorem 1.6 both achieved, i.e.,
-$\min_{xy\in E}\kappa(x,y)=\lambda_{2}=\frac{2+\alpha}{d},$
-have been fully classified in [14, Theorem 6.4]. Accordingly, we obtain the following classification of Licherowicz sharp strongly regular graphs with $\beta\geq\alpha$.
-Corollary 1.8.
-The connected Lichnerowicz sharp strongly regular graphs with parameters $(n,d,\alpha,\beta)$ such that $\beta\geq\alpha$ are precisely the following ones: The cocktail party graphs $CP(k),k\geq 2$, the Paley graph $P(9)$, the Johnson graphs $J(5,2)$ and $J(6,2)$ and the demi-cube $Q^{5}_{(2)}$ (also called Clebsch graph).
-The rest of the our paper is organized as follows.
+Consider two elements $x,y\in GF(n)$ such that xyx-y is a non-zero square in $GF(n)$.
+For any subset $S\subset GF(n)\setminus\{x,y\}$ with $|S|\geq\frac{3}{4}(n-1)$, there exist $w,z\in S$ such that xw,wz,zyx-w,w-z,z-y are non-zero squares and xz,ywx-z,y-w are not squares in $GF(n)$.
+In particular, the above corollary applies to quadratic residues when $n$ is a Pythagorean prime (i.e., a prime congruent to $1$ modulo $4$) greater than $5$.
+The rest of this paper is organized as follows.
 In Section 2, we recall definitions and properties about the Lin–Lu–Yau curvature, perfect matching, and strongly/amply regular graphs.
-In Section 3, we prove Theorem 1.3, Theorem 1.4 and Corollary 1.5, and present possible ways to generalize our conclusion.
-In Section 4, we prove the Theorem 1.7 and Corollary 1.8.
+In Section 3, we prove Theorem 1.3, Theorem 1.4, and Corollary 1.5, and present possible ways to generalize our conclusion.
 2 Preliminaries
@@
 where $d(x,y)$ denotes the combinatorial distance between $x$ and $y$ in $G$, and the infimum is taken over all maps $\pi:V\times V\to[0,1]$ satisfying
-$\mu_{1}(x)=\sum_{y\in V}\pi(x,y),\,\,\mu_{2}(y)=\sum_{x\in V}\pi(x,y).$
+[…]
 Such a map is called a transport plan.
 We consider the following particular measure around a vertex $x\in V$:
-$\mu_{x}^{p}(y)=\left\{\begin{array}[]{ll}p,&\hbox{if $y=x$;}\\ \frac{1-p}{\mathrm{deg}(x)},&\hbox{if $y\sim x$;}\\ 0,&\hbox{otherwise,}\end{array}\right.$
-where $\mathrm{deg}(x):=\sum_{y\in V:y\sim x}1$ is the vertex degree of $x$.
+$\mu_{x}^{p}(y)=\left\{\begin{array}[]{ll}p,&\hbox{if $y=x$;}\\ \frac{1-p}{\mathrm{deg}(x)},&\hbox{if $yx\in E$;}\\ 0,&\hbox{otherwise,}\end{array}\right$.
+where $\mathrm{deg}(x)$ is the degree of the vertex $x$.
 Definition 2.2 ($p$-Ollivier curvature [26] and Lin–Lu–Yau curvature [21]).
@@
 $\kappa(x,y)=\lim_{p\to 1}\frac{\kappa_{p}(x,y)}{1-p}.$
-Notice that $\kappa_{1}(x,y)$ is always 00. Hence, the Lin–Lu–Yau curvature $\kappa(x,y)$ equals the negative of left derivative of the function $p\mapsto\kappa_{p}(x,y)$ at $p=1$.
-Boure et al. [7] proved the following relation between $p$-Ollivier curvature and Lin–Lu–Yau curvature for an edge $xy\in E$ of a $d$-regular graph $G=(V,E)$:
+Notice that $\kappa_{1}(x,y)$ is always 0. Hence, the Lin–Lu–Yau curvature $\kappa(x,y)$ is equal to the negative of the left derivative of the function $p\mapsto\kappa_{p}(x,y)$ at $p=1$.
+Bourne et al. [7] proved the following relation between $p$-Ollivier curvature and Lin–Lu–Yau curvature for an edge $xy$ of a $d$-regular graph $G$:
 $\kappa(x,y)=\frac{d+1}{d}\kappa_{\frac{1}{d+1}}(x,y).$ (2.1)
-This relation leads in particular to the following expression of Lin–Lu–Yau curvature of a $d$-regular graph:
+By the property of the Monge problem, this relation leads in particular to the following expression of Lin–Lu–Yau curvature of a $d$-regular graph:
 $\kappa(x,y)=\frac{1}{d}\left(d+1-\min_{\begin{subarray}{c}\phi:N_{x}\to N_{y}\\ \text{bijective}\end{subarray}}\sum_{v\in N_{x}}d(v,\phi(v))\right).$ (2.2)
@@
 A matching $M$ is called a perfect matching if it covers every vertex of the graph.
-Hall’s marriage Theorem is an fundamental tool for proving Conjecture 1.2.
+Hall’s marriage Theorem is a fundamental tool for proving Conjecture 1.2.
 Theorem 2.4.
@@
 $|V_{1}|=|V_{2}|\,\,\text{and}\,\,|\Gamma_{V_{2}}(S)|\geq|S|\,\,\text{for all}\,\,S\subseteq V_{1},$
-holds, where $\Gamma_{V_{2}}(S):=\{v\in V_{2}|\,\,\text{there exists }\,w\in S\,\text{such that}\,\,vw\in E\}$.
+where $\Gamma_{V_{2}}(S):=\{v\in V_{2}|\,\,\text{there exists }\,w\in S\,\text{such that}\,\,vw\in E\}$.
 2.2 Strongly/Amply regular graph
 Definition 2.5 (Strongly/Amply regular graph [8]).
-Let $G$ be a $d$-regular graph with $n$ vertices which is neither complete nor empty. Then $G$ is called call a strongly regular graph with parameters $(n,d,\alpha,\beta)$ if:
+Let $G$ be a $d$-regular graph with $n$ vertices which is neither complete nor empty.
+The graph $G$ is called a strongly regular graph with parameters $(n,d,\alpha,\beta)$ if:
   • (i)
@@
     Any two non-adjacent vertices have $\beta$ common neighbors.
-[…]
+The graph $G$ is called an amply regular graph with parameters $(n,d,$ $\alpha,\beta)$ if:
   • (i)
@@
   • (ii)
-    Any two vertices with distance $2$ have $\beta$ common neighbors.
-The following theorem shows a relationship between the parameters of an amply regular graph.
+    Any two vertices at distance two have $\beta$ common neighbors.
+The following theorem shows a basic relationship between the parameters of an amply regular graph.
 Theorem 2.6.
-Let $G=(V,E)$ be an amply regular graph with parameters $(n,d,\alpha,\beta)$. then
+Let $G=(V,E)$ be an amply regular graph with parameters $(n,d,\alpha,\beta)$. Then,
 $d(d-\alpha-1)\leq(n-d-1)\beta.$ (2.3)
@@
 Proof.
-Let $M$ be the number of pairs of ($(\{v_{1},v_{2}\},u)$, where $v_{1},v_{2},u\in V$ and $u$ is a common neighbour of $v_{1}$ and $v_{2}$.
+Let $M$ be the number of pairs of $(\{v_{1},v_{2}\},u)$, where $v_{1},v_{2},u\in V$ and $u$ is a common neighbour of $v_{1}$ and $v_{2}$.
 For any $u\in V$, there are $\binom{d}{2}$ pairs of such $\{v_{1},v_{2}\}$.
@@
 For any two adjacent vertices $v_{1}$ and $v_{2}$, there are $\alpha$ such $u$.
-For any two non-adjacent vertices $v_{1}$ and $v_{2}$, there are at either $\beta$ or 00 such $u$.
+For any two non-adjacent vertices $v_{1}$ and $v_{2}$, there are at either $\beta$ or 0 such $u$.
 Therefore,
 $M\leq e\alpha+\left(\binom{n}{2}-e\right)\beta,$
-where $e=|E|=\frac{nd}{2}$, and the inequality (2.3) follows. It’s clear that the equality holds if and only if $G$ is strongly regular. ∎
-3 Proof of Theorem 1.3, Theorem 1.4 and Corollary 1.5
-In this section, we prove Theorem 1.3, Theorem 1.4 and Corollary 1.5. First, we observe the following lemma which is useful in applying Hall’s marriage Theorem (Theorem 2.4).
+where $e=|E|=\frac{nd}{2}$, and the inequality (2.3) follows. It is clear that the equality holds if and only if $G$ is strongly regular. ∎
+3 Proof of Theorem 1.3, Theorem 1.4, and Corollary 1.5
+In this section, we prove Theorem 1.3, Theorem 1.4, and Corollary 1.5. First, we observe the following lemma which is useful in applying Hall’s marriage Theorem (Theorem 2.4).
 Lemma 3.1.
@@
 Lemma 3.2.
-Let $G=(V,E)$ be an amply regular graphs with parameters $(n,d,\alpha,\beta)$ such that $d\geq\alpha+2$. We have the following properties:
-  • (i)
-    If $n<3d-2\alpha$, then $d\leq 2\beta$.
-  • (ii)
-    If $2n<5d-3\alpha$, then $d\leq\frac{3}{2}\beta$.
+Let $G=(V,E)$ be an amply regular graphs with parameters $(n,d,\alpha,\beta)$ such that $d\geq\alpha+2$. If $n<3d-2\alpha$, then $d\leq 2\beta$.
 Proof.
@@
 $d(d-\alpha-1)\leq(n-d-1)\beta\leq 2(d-\alpha-1)\beta.$ (3.1)
-Since $d\geq\alpha+2$, we deduce $d\leq 2\beta$. This proves $(i)$. $(ii)$ is deduced similarly. ∎
+Since $d\geq\alpha+2$, we deduce $d\leq 2\beta$. ∎
 Let $x,y$ be two adjacent vertices of an amply regular graph $G=(V,E)$ with parameters $(n,d,\alpha,\beta)$.
@@
 Figure 1: A schematic plot for the decomposition of $V$
-The subsets above are defined by $\Delta_{xy}:=\Gamma(x)\cap\Gamma(y)$, $N_{x}:=\Gamma(x)\backslash(\Delta_{xy}\cup\{y\})$, $N_{y}:=\Gamma(y)\backslash(\Delta_{xy}\cup\{x\})$, and $P_{xy}:=V\backslash(\Gamma(x)\cup\Gamma(y))$, where $\Gamma(x):=\{z\in V|\,xz\in E\}$, and $\Gamma(y):=\{z\in V|\,yz\in E\}$.
+The subsets above are defined by $\Delta_{xy}:=\Gamma(x)\cap\Gamma(y)$, $N_{x}:=\Gamma(x)\backslash(\Delta_{xy}\cup\{y\})$, $N_{y}:=\Gamma(y)\backslash(\Delta_{xy}\cup\{x\})$, and $P_{xy}:=V\backslash(\Gamma(x)\cup\Gamma(y))$, where $\Gamma(x):=\{z\in V|\,xz\in E\}$ and $\Gamma(y):=\{z\in V|\,yz\in E\}$.
 It is direct to check by definition that
@@
 Since $x$ and $v$ have $\alpha$ common neighbors, $v$ has $\alpha-\beta+1+\ell$ neighbors in $N_{x}$.
-Since $v$ has $d-1$ neighbors except $x$, the number of neighbors of $v$ in $P_{xy}$ is $d-\alpha-1-\ell$.
+Since $v$ has d1d-1 neighbors except $x$, the number of neighbors of $v$ in $P_{xy}$ is $d-\alpha-1$-$\ell$.
 ∎
-Now, we are prepared to prove Theorem 1.4.
+We now proceed to the proof of Theorem 1.3.
 Proof of Theorem 1.3.
-Let $G=(V,E)$ be an amply regular graph satisfying one of the conditions (1)–(5), and $x,y\in V$ be two adjacent vertices.
-Recall from [19, Proposition 3.6] that $\kappa(x,y)=\frac{2+\alpha}{d}$ holds whenever $d\leq 2\beta-\alpha-1$.
-In the following proof, we assume that $d\geq 2\beta-\alpha$.
+Let $G=(V,E)$ be an amply regular graph satisfying one of the conditions (1)–(3), and let $xy\in E$ be an edge.
+Recall from [17, Proposition 3.6] that $\kappa(x,y)=\frac{2+\alpha}{d}$ holds whenever $d\leq 2\beta-\alpha-1$.
+Thus, we assume throughout the proof that $d\geq 2\beta-\alpha$.
 We claim that each of the conditions (1)–(3) yields $d>\alpha+2$.
-Indeed, it is clear that conditions (1) and (2) yield $d>\alpha+2$.
-Under the condition (3) or (4), $d\geq\beta>\alpha+2$.
-Under the condition (5), $d\geq 2\alpha+3-\frac{(\beta-1)^{2}}{\alpha}>\alpha+3$.
+Indeed, it is clear that condition (1) yields $d>\alpha+2$.
+Under the condition (2), $d\geq\frac{3}{2}\beta>\alpha+2$.
+Since $d\geq\beta$, condition (3) implies $d>\alpha+2$.
 Therefore, the claim holds.
@@
 That is, both $N_{x}$ and $N_{y}$ contain at least two vertices.
-Recall from [22, Theorem 1.5] that $\kappa(x,y)=\frac{2+\alpha}{d}$ holds whenever $\alpha=0$ and $\beta\geq 2$.
-If $\alpha=0$, then each of the conditions (2)–(4) leads to $\beta\geq 2$ and hence $\kappa(x,y)=\frac{2}{d}$.
+Recall from [20, Theorem 1.5] that $\kappa(x,y)=\frac{2+\alpha}{d}$ holds whenever $\alpha=0$ and $\beta\geq 2$.
+If $\alpha=0$, then the conditions (2) and (3) both lead to $\beta\geq 2$ and hence $\kappa(x,y)=\frac{2}{d}$.
 Under the condition (1), $\alpha=0$ implies $\beta=1$ and $d>2$.
@@
 That is, under the condition (1), the case $\alpha=0$ is not allowed.
-Under the condition (5), it’s clear that the case $\alpha=0$ can not happen.
 In the following proof, we assume that $\alpha>0$.
@@
 always holds.
-For each vertex $v_{i}\in S:=\{v_{1},v_{2},\ldots,v_{b}\}$, let $x_{i}$ be the number of neighbors of $v_{i}$ in $N_{y}$.
+For each vertex $v_{i}\in S$, let $x_{i}$ be the number of neighbors of $v_{i}$ in $N_{y}$.
 By Lemma 3.3, the vertex $v_{i}$ has $d-\alpha-1-x_{i}$ neighbors in $P_{xy}$.
@@
 By symmetry and Lemma 3.1, we only need to show the inequality (3.3) for $b\leq\frac{d-\alpha}{2}$.
-We argue by contradiction. Suppose $|T|\leq b-1$.
+Assume, for the sake of contradiction, that $|T|\leq b-1$.
 Case 1: $b=1$.
-If $b=1$, then $v_{1}$ has no neighbor in $N_{y}$, that is, $x_{1}=0$.
-Thus by Lemma 3.3, $v_{1}$ has $\beta-1$ neighbors in $\Delta_{xy}$ and $d-\alpha-1$ neighbors in $P_{xy}$.
-If condition (2) or (3) or (4) holds, then $\beta-1>\alpha=|\Delta_{xy}|$, which is a contradiction.
-If condition (5) holds, then $|P_{xy}|=n-2d+\alpha\leq\frac{1}{2}(d-\alpha-1)<d-\alpha-1$, which is a contradiction.
+If $b=1$, then $v_{1}$ has no neighbors in $N_{y}$, that is, $x_{1}=0$.
+Thus, by Lemma 3.3, $v_{1}$ has $\beta-1$ neighbors in $\Delta_{xy}$ and $d-\alpha-1$ neighbors in $P_{xy}$.
+If condition (2) or (3) holds, then $\beta-1>\alpha=|\Delta_{xy}|$, which is a contradiction.
 If condition (1) holds, we have $\beta-1=\alpha$ and hence $v_{1}$ is adjacent to all vertices in $\Delta_{xy}$.
@@
 Thus, $|P_{xy}|=d-\alpha-1$ and $v_{1}$ is adjacent to all vertices in $P_{xy}$.
-Therefore, $v_{1}$ has in total $d-1$ neighbors in $\Delta_{xy}\cup P_{xy}$.
+Therefore, $v_{1}$ has in total d1d-1 neighbors in $\Delta_{xy}\cup P_{xy}$.
 Hence, $v_{1}$ has no neighbors in $N_{x}\cup N_{y}$.
@@
 that is,
-$2\displaystyle 2q=d-\alpha-$ 2 . (3.5)
-Thus, $d-\alpha-1$ is odd, which is a contradiction.
+$2q=d-\alpha-2.$ (3.5)
+Thus, $d-\alpha-1$ is odd, contradicting (3.4).
 Case 2: $2\leq b\leq\frac{d-\alpha}{2}$.
-Let $X=x_{1}+\cdots+x_{b}$, then $X$ is the number of edges between $S:=\{v_{1},\ldots,v_{b}\}$ and $T$.
+Let $X=x_{1}+\cdots+x_{b}$, then $X$ is the number of edges between $S$ and $T$.
 Set $\Delta_{xy}:=\{\delta_{1},\ldots,\delta_{\alpha}\}$, and let $A_{k}$ be the number of neighbors of $\delta_{k}$ in $S$, where $1\leq k\leq\alpha$.
@@
 $M_{2}:=\sum_{1\leq i<j\leq b}Y_{ij}.$
-Since $n-2d+\alpha\leq d-\alpha-1$ under each condition $(1)-(5)$, it follows by Cauchy inequality that
+Since $n-2d+\alpha\leq d-\alpha-1$ under each condition (1)(5)(1)-(5), it follows by Cauchy inequality that
 $M_{2}=\sum_{k=1}^{n-2d+\alpha}\binom{B_{k}}{2}$ $\geq\frac{1}{2}\left[\frac{1}{n-2d+\alpha}\left(\sum_{k=1}^{n-2d+\alpha}B_{k}\right)^{2}-\sum_{k=1}^{n-2d+\alpha}B_{k}\right]$
@@
 $M_{3}\geq\frac{1}{2}\left[\frac{1}{|T|}\left(\sum_{k=1}^{|T|}C_{k}\right)^{2}-\sum_{k=1}^{|T|}C_{k}\right]\geq\frac{1}{2}\left[\frac{1}{b-1}X^{2}-X\right].$ (3.9)
-Set $N_{x}=\{v_{1},\ldots,v_{d-\alpha-1}\}$, and let $D_{k}$ be the number of neighbors of $v_{k}$ in $S$, where $11\leq k\leq d-\alpha-$ 1.
+Set $N_{x}=\{v_{1},\ldots,v_{d-\alpha-1}\}$, and let $D_{k}$ be the number of neighbors of $v_{k}$ in $S$, where $1\leq k\leq d-\alpha-1$.
 Counting the number of ordered pairs $(v_{k},v_{j})$ with $v_{k}\in N_{x}$, $v_{j}\in S$ and $v_{k}v_{j}\in E$, we have
@@
 $\leq\binom{b}{2}\max\{\alpha,\beta\}.$ (3.11)
-We divide the remaining argument of Case 2 into four subcases.
+We divide the remaining argument of Case 2 into two subcases.
 (i) Suppose that the condition (1) holds. Inserting (3), (3.8), (3.9), (3) into (3) yields
@@
 which is a contradiction.
-(ii) Suppose that the condition (2) holds. Then we derive from (3) that
-$M_{1}\geq\frac{1}{2}\left[\frac{\big{(}(\alpha+1)b-X\big{)}^{2}}{\alpha+1}-\big{(}(\alpha+1)b-X\big{)}\right].$ (3.14)
-Inserting (3.14), (3.8), (3.9), and (3) into (3), we have
-$\left(\frac{1}{\alpha+1}+\frac{2}{d-\alpha-1}+\frac{1}{b-1}\right)X^{2}$ $-\left(4b+\frac{2b}{d-\alpha-1}\right)X$
-$+\frac{b^{2}}{d-\alpha-1}+b+(b^{2}-b)(d-\alpha-1)\leq 0.$
-Since the discriminant of the above quadratic polynomial is non-negative, we deduce that
-$b\left(2+\frac{1}{d-\alpha-1}\right)^{2}$ $\geq\left[\frac{1}{\alpha+1}+\frac{2}{d-\alpha-1}+\frac{1}{b-1}\right]\left[\frac{b}{d-\alpha-1}+1+(b-1)(d-\alpha-1)\right]$
-$>2(b-1)+d-\alpha-1+\frac{b}{(d-\alpha-1)^{2}}+\frac{2}{d-\alpha-1}+\frac{(b-1)(d-\alpha-1)}{\alpha+1}.$
-That is,
-$0>d-\alpha-2+\frac{2}{d-\alpha-1}-\frac{d}{\alpha+1}-\left(3+\frac{4}{d-\alpha-1}-\frac{d}{\alpha+1}\right)b.$ (3.15)
-Since $d\leq 2\beta$, we obtain
-$3+\frac{4}{d-\alpha-1}-\frac{d}{\alpha+1}>3-\frac{d}{\alpha+1}\geq 3-\frac{2\beta}{\alpha+1}\geq 0,$
-and hence
-00 0 $>d-\alpha-2+\frac{2}{d-\alpha-1}-\frac{d}{\alpha+1}-\left(3+\frac{4}{d-\alpha-1}-\frac{d}{\alpha+1}\right)\frac{d-\alpha}{2}$
-$=d-\alpha-4-\frac{3(d-\alpha)}{2}+\frac{d(d-\alpha-2)}{2(\alpha+1)}.$ (3.16)
-Considering the above inequality as a quadratic inequality with respect to $d$, we deduce
-$d<\alpha+\sqrt{10\alpha+\frac{41}{4}}+\frac{3}{2},$
-which is a contradiction.
-(iii) Suppose that either condition (3) or (4) holds. Inserting (3), (3.8), (3.9) and $M_{4}\geq 0$ into the inequality (3) leads to
+(ii) Suppose that either condition (2) or (3) holds. Inserting (3), (3.8), (3.9) and $M_{4}\geq 0$ into the inequality (3) leads to
 $\left(\frac{1}{\alpha}+\frac{1}{d-\alpha-1}+\frac{1}{b-1}\right)X^{2}$ $-\left(\frac{2b(\beta-1)}{\alpha}+2b-1\right)X$
@@
 Using $\frac{2b(\beta-1)}{\alpha}+2b-1<\frac{2b(\beta-1)}{\alpha}+2b$, we derive from the non-negativity of the discriminant of the above quadratic polynomial that
-$b\left(\frac{\beta-1}{\alpha}+1\right)^{2}>\left[\frac{1}{\alpha}+\frac{1}{d-\alpha-1}+\frac{1}{b-1}\right]\left[\frac{b(\beta-1)^{2}}{\alpha}-(\beta-1)+(b-1)(d-\alpha-\beta)\right].$
+$b\left(\frac{\beta-1}{\alpha}+1\right)^{2}\geq\left[\frac{1}{\alpha}+\frac{1}{d-\alpha-1}+\frac{1}{b-1}\right]\left[\frac{b(\beta-1)^{2}}{\alpha}-(\beta-1)+(b-1)(d-\alpha-\beta)\right].$
 That is,
-$\frac{2b(\beta-1)}{\alpha}+2b>$ $1+\frac{bd-b\beta-d+1}{\alpha}$
-$+\left[\frac{1}{d-\alpha-1}+\frac{1}{b-1}\right]\left[\frac{b(\beta-1)^{2}}{\alpha}-(\beta-1)+(b-1)(d-\alpha-\beta)\right].$
+$\frac{2b(\beta-1)}{\alpha}+2b$
+$\geq$ $1+\frac{bd-b\beta-d+1}{\alpha}+\left[\frac{1}{d-\alpha-1}+\frac{1}{b-1}\right]\left[\frac{b(\beta-1)^{2}}{\alpha}-(\beta-1)+(b-1)(d-\alpha-\beta)\right].$
 Since
@@
 $\frac{2b(\beta-1)}{\alpha}+b$
-$>$ $\frac{bd-b\beta-d+1}{\alpha}+d-\alpha-1+\left(\frac{1}{d-\alpha-1}+\frac{1}{b-1}\right)b(\beta-1)\left(\frac{\beta-1}{\alpha}-1\right)$
-$>$ $\frac{bd-b\beta-d+1}{\alpha}+d-\alpha-1+(\beta-1)\left(\frac{\beta-1}{\alpha}-1\right)+\frac{b(\beta-1)}{d-\alpha-1}\left(\frac{\beta-1}{\alpha}-1\right).$
+$\geq$ $\frac{bd-b\beta-d+1}{\alpha}+d-\alpha-1+\left(\frac{1}{d-\alpha-1}+\frac{1}{b-1}\right)b(\beta-1)\left(\frac{\beta-1}{\alpha}-1\right)$
+$>$ $\frac{bd-b\beta-d+1}{\alpha}+d-\alpha-1+\frac{b(\beta-1)}{d-\alpha-1}\left(\frac{\beta-1}{\alpha}-1\right)+(\beta-1)\left(\frac{\beta-1}{\alpha}-1\right).$
 That is,
 $\frac{-d+1}{\alpha}+d-\alpha-1+(\beta-1)\left(\frac{\beta-1}{\alpha}-1\right)$
-$<\left[\frac{2(\beta-1)}{\alpha}+1-\frac{d-\beta}{\alpha}-\frac{\beta-1}{d-\alpha-1}\left(\frac{\beta-1}{\alpha}-1\right)\right]b.$
+$<$ $\left[\frac{2(\beta-1)}{\alpha}+1-\frac{d-\beta}{\alpha}-\frac{\beta-1}{d-\alpha-1}\left(\frac{\beta-1}{\alpha}-1\right)\right]b.$ (3.14)
 Since $2\beta-\alpha\leq d\leq 2\beta\leq\alpha+\frac{5}{2}\beta-\frac{3}{2}$, we have
@@
 $\geq$ $\frac{2(\beta-1)}{\alpha}-\frac{3(\beta-1)}{2\alpha}-\frac{\beta-1}{2(\beta-\alpha-1)}\left(\frac{\beta-1}{\alpha}-1\right)=0.$
-Therefore, we obtain
+Therefore, by the assumption that $b\leq\frac{d-\alpha}{2}$, the inequality (3.14) yields
 $\frac{-d+1}{\alpha}+d-\alpha-1+(\beta-1)\left(\frac{\beta-1}{\alpha}-1\right)$
@@
 By (3.15), we have
-$\left|d-\frac{3}{2}\beta\right|<\frac{1}{2}\sqrt{4\alpha^{2}-3\beta^{2}+4\alpha+24\beta-20}).$
+$\left|d-\frac{3}{2}\beta\right|<\frac{1}{2}\sqrt{4\alpha^{2}-3\beta^{2}+4\alpha+24\beta-20})<\frac{1}{2}\sqrt{4\alpha^{2}-3\beta^{2}+28\beta}).$
 Therefore, neither condition (2) nor (3) holds, which is a contradiction.
-(iv) Suppose that the condition (5) holds. Then we have $n-2d+\alpha\leq\frac{1}{2}(d-\alpha-1)$. It follows from (3.7) that
-$M_{2}\geq\frac{1}{2}\left[\frac{2\left((d-\alpha-1)b-X\right)^{2}}{d-\alpha-1}-\left((d-\alpha-1)b-X\right)\right].$ (3.18)
-Inserting (3), (3.18), (3.9) and (3) into the inequality (3), we have
-$\left(\frac{1}{\alpha}+\frac{3}{d-\alpha-1}+\frac{1}{b-1}\right)X^{2}-\left(\frac{2b(\beta-1)}{\alpha}+4b-\frac{2(\alpha-\beta+1)b}{d-\alpha-1}\right)X$
-$\hskip 19.91684pt+\frac{b^{2}(\beta-1)^{2}}{\alpha}+2(d-\alpha-1)b^{2}+\frac{(\alpha-\beta+1)^{2}b^{2}}{d-\alpha-1}-(d-\alpha)b-(\alpha-1)b^{2}\leq 0.$ (3.19)
-Since the discriminant of the above quadratic polynomial is non-negative, we deduce
-$b\left(\frac{\beta-1}{\alpha}+2-\frac{\alpha-\beta+1}{d-\alpha-1}\right)^{2}\geq\left(\frac{1}{\alpha}+\frac{3}{d-\alpha-1}+\frac{1}{b-1}\right)\times$
-$\hskip 19.91684pt\left(\frac{b(\beta-1)^{2}}{\alpha}+2(d-\alpha-1)b+\frac{(\alpha-\beta+1)^{2}b}{d-\alpha-1}-(d-\alpha)-(\alpha-1)b\right).$ (3.20)
-It is clear that
-$\frac{1}{\alpha}\left(\frac{b(\beta-1)^{2}}{\alpha}+2(d-\alpha-1)b+\frac{(\alpha-\beta+1)^{2}b}{d-\alpha-1}-(d-\alpha)-(\alpha-1)b\right)$
-$>$ $\frac{b(\beta-1)^{2}}{\alpha^{2}}+\frac{2b(d-\alpha-1)}{\alpha}-\frac{d-\alpha+\alpha b-b}{\alpha}.$ (3.21)
-Since $d\geq 2\alpha+3-\frac{(\beta-1)^{2}}{\alpha}$, we have
-$d-\alpha-1\geq\alpha+2-\frac{(\beta-1)^{2}}{\alpha}>\alpha-\beta+3\geq 3.$ (3.22)
-Thus, we estimate
-$\frac{3}{d-\alpha-1}\left(\frac{b(\beta-1)^{2}}{\alpha}+2(d-\alpha-1)b+\frac{(\alpha-\beta+1)^{2}b}{d-\alpha-1}-(d-\alpha)-(\alpha-1)b\right)$
-$=$ $\frac{3b(\beta-1)^{2}}{(d-\alpha-1)\alpha}+6b+\frac{3(\alpha-\beta+1)^{2}b}{(d-\alpha-1)^{2}}-3-\frac{3}{d-\alpha-1}-\frac{3(\alpha-1)b}{d-\alpha-1}$
-$>$ $\frac{3b(\beta-1)^{2}}{(d-\alpha-1)\alpha}+6b+\frac{(\alpha-\beta+1)^{2}b}{(d-\alpha-1)^{2}}-4-\frac{3(\alpha-1)b}{d-\alpha-1}.$ (3.23)
-By the condition that $d\geq 2\alpha+3-\frac{(\beta-1)^{2}}{\alpha}$, we derive
-$\frac{1}{b-1}\left(\frac{b(\beta-1)^{2}}{\alpha}+2(d-\alpha-1)b+\frac{(\alpha-\beta+1)^{2}b}{d-\alpha-1}-(d-\alpha)-(\alpha-1)b\right)$
-$=$ $\frac{(\beta-1)^{2}}{\alpha}+2(d-\alpha-1)+\frac{(\alpha-\beta+1)^{2}}{d-\alpha-1}-(\alpha-1)$
-$\hskip 39.83368pt+\frac{1}{b-1}\left(\frac{(\beta-1)^{2}}{\alpha}+2(d-\alpha-1)+\frac{(\alpha-\beta+1)^{2}}{d-\alpha-1}-(d-1)\right)$
-$>$ $\frac{(\beta-1)^{2}}{\alpha}+2(d-\alpha-1)+\frac{(\alpha-\beta+1)^{2}}{d-\alpha-1}-(\alpha-1)$
-$>$ $\frac{(\beta-1)^{2}}{\alpha}+2(d-\alpha-1)+\frac{2(\alpha-\beta+1)^{2}}{\alpha}-(\alpha-1),$ (3.24)
-where the last inequality above follows from Lemma 3.2. Inserting (3.21),(3),(3) into (3.20) yields
-$-\frac{d}{\alpha}+\frac{(\beta-1)^{2}}{\alpha}+2(d-\alpha-1)+\frac{2(\alpha-\beta+1)^{2}}{\alpha}-\alpha-2$
-$<$ $\left(\frac{4\beta-2d-3}{\alpha}-\frac{\alpha-\beta+4}{d-\alpha-1}+\frac{(\beta-1)(\alpha-\beta+1)}{\alpha(d-\alpha-1)}+1\right)b.$ (3.25)
-By Lemma 3.2, we have $\frac{3}{2}\beta\geq d$. By (3.22), we have $d-\alpha-1\geq\alpha-\beta+4$. Thus, the coefficient of $b$ satisfies
-$\frac{4\beta-2d-3}{\alpha}-\frac{\alpha-\beta+4}{d-\alpha-1}+\frac{(\beta-1)(\alpha-\beta+1)}{\alpha(d-\alpha-1)}+1\geq\frac{\beta-3}{\alpha}-1+0+1\geq 0.$
-It follows from (3) that
-$-\frac{d}{\alpha}+\frac{(\beta-1)^{2}}{\alpha}+2(d-\alpha-1)+\frac{2(\alpha-\beta+1)^{2}}{\alpha}-\alpha-2$
-$<$ $\left(\frac{4\beta-2d-3}{\alpha}-\frac{\alpha-\beta+4}{d-\alpha-1}+\frac{(\beta-1)(\alpha-\beta+1)}{\alpha(d-\alpha-1)}+1\right)\frac{d-\alpha}{2}$
-$\leq$ $\frac{(4\beta-2d-3)(d-\alpha)}{2\alpha}-\frac{\alpha-\beta+4}{2}+\frac{(\beta-1)(\alpha-\beta+1)}{2\alpha}+\frac{(\beta-1)(\alpha-\beta+1)}{2\alpha(d-\alpha-1)}+\frac{d-\alpha}{2}$
-$<$ $\frac{(4\beta-2d-3)(d-\alpha)}{2\alpha}-\frac{\alpha-\beta+4}{2}+\frac{(\beta-1)(\alpha-\beta+1)}{2\alpha}+\frac{1}{2}+\frac{d-\alpha}{2}.$ (3.26)
-That is,
-$02d^{2}+(\alpha-4\beta+1)d-6\alpha\beta+\alpha+7\beta^{2}-14\beta+7<$ 0 .
-Since $06\beta-d-1>$ 0, we have
-$\alpha>\frac{2d^{2}-4\beta d+7\beta^{2}+d-14\beta+7}{6\beta-d-1},$
-which is a contradiction.
 Now, we have completed Case 2 and established Theorem 1.3.∎
 Remark 3.4.
-In addition to the cases in Theorem 1.4, our method can deal with many other cases. Here, we comment on some possible ways to extend our conclusion.
+In addition to the cases in Theorem 1.3 and [9, Theorem 1.4], our method can deal with many other cases. Here, we comment on some possible ways to extend our conclusion.
   • (i)
@@
   • (ii)
-    Suppose that there are $e$ edges within $S:=\{v_{1},\ldots,v_{b}\}$ and $f$ edges between $S$ and $N_{x}\backslash S$.
-Note that $\sum_{k=1}^{b}D_{k}=2e$ and $\sum_{k=b+1}^{d-\alpha-1}D_{k}=f$.
-The inequality (3) can be improved as
+[…]
     $M_{4}=\sum_{k=1}^{d-\alpha-1}\binom{D_{k}}{2}\geq\frac{1}{2}\biggl(\frac{(2e)^{2}}{b}+\frac{f^{2}}{d-\alpha-1-b}-\sum_{k=1}^{d-\alpha-1}D_{k}\biggr),$
@@
     These modifications might improve our conclusion to some extent.
-To illustrate the above remark, we show that the Lin–Lu–Yau curvature of a particular strongly regular graph with $n\geq 3d-2\alpha$ achieves the upper bound $\frac{2+\alpha}{d}$.
+To illustrate the above remark (i), we show that the Lin–Lu–Yau curvature of a particular strongly regular graph with $n\geq 3d-2\alpha$ achieves the upper bound $\frac{2+\alpha}{d}$.
 Example 3.5.
-Consider a strongly regular graph $G$ with parameters (324, 152, 70, 72) derived from regular symmetric Hadamard matrices with constant diagonal (RSHCD), which is constructed in [11, Section 5.3.2] based on [18, Lemma 11].
+Consider a strongly regular graph $G$ with parameters (324, 152, 70, 72) derived from regular symmetric Hadamard matrices with constant diagonal (RSHCD), which is constructed in [11, Section 5.3.2] based on [16, Lemma 11].
 Note that the condition $n<3d-2\alpha$ does not hold.
@@
 $\left(\frac{107}{5670}+\frac{1}{2(b-1)}\right)X^{2}-\frac{5462b}{2835}X+\frac{104791b^{2}}{2835}-40b\leq 0,$
-where $412\leq b\leq\frac{d-\alpha}{2}=$ 41. Considering the discriminant, we have
+where $2\leq b\leq\frac{d-\alpha}{2}=41$. Considering the discriminant, we have
 $\left(\frac{5462b}{2835}\right)^{2}\geq 4\left(\frac{107}{5670}+\frac{1}{b-1}\right)\left(\frac{104791b^{2}}{2835}-40b\right).$
-It is direct to check that the inequality above can not hold for $2\leq b\leq 41$. Therefore, the Lin–Lu–Yau curvature of $G$ is
+It is direct to check that the inequality above does not hold for $2\leq b\leq 41$. Therefore, the Lin–Lu–Yau curvature of $G$ is
 $\frac{2+\alpha}{d}=\frac{9}{19}.$
-Now, let’s prove Theorem 1.3.
+Now, let us prove Theorem 1.4.
 Recall that, in the setting of Theorem 1.4, we have $(n,d,\alpha,\beta)=(4\gamma+1,2\gamma,\gamma-1,\gamma)$.
-By the condition $(1)$ of Theorem 1.4, we only need to verify the cases that $\gamma$ is equal to $2$, $3$, $4$, $5$, or $6$.
+By the condition $(1)$ of Theorem 1.3, we only need to verify the cases that $\gamma$ is equal to $2$, 33, 44, 55, or $6$.
 Proof of Theorem 1.4.
@@
 Suppose that $\gamma=5$, then $b=3$. Substituting $\gamma=5$ and $b=3$ into the inequality (3.12), we have
-$30-12X+\frac{23X^{2}}{20}\leq 0$
+$30-12X+\frac{23X^{2}}{20}\leq 0.$
 It follows that $X=5$ or $X=6$.
-In case that $X=5$, we derive from (3), (3.8) and (3.9) that $M_{1}\geq 3$, $M_{2}\geq 5$ and $M_{3}\geq 4$ respectively.
-We claim that $M_{4}\geq 1$.
-Set $N_{x}=\{v_{1},\ldots,v_{5}\}$.
-Denote by $e$ the number of edges within $\{v_{1},\ldots,v_{5}\}$ and $f$ edges between $\{v_{1},\ldots,v_{5}\}$ and $N_{x}\backslash\{v_{1},\ldots,v_{5}\}$.
-Observe that $X=2e+f$ by Lemma 3.3 (ii).
-Thus, the number of edges within $\{v_{1},v_{2},v_{3}\}$ is at most $2$.
-If there are 2 edges within $\{v_{1},v_{2},v_{3}\}$, then $M_{4}\geq 1$, as claimed.
-If the number of edges within $\{v_{1},v_{2},v_{3}\}$ is at most $1$, then there are at least $3$ edges between $\{v_{1},v_{2},v_{3}\}$ and $\{v_{4},v_{5}\}$, which implies that either $v_{4}$ or $v_{5}$ is adjacent to two vertices of $\{v_{1},v_{2},v_{3}\}$.
-That is, we have $M_{4}\geq 1$ as claimed.
+[…]
 In case that $X=6$, we derive from (3), (3.8), (3.9) and (3) that $M_{1}\geq 2$, $M_{2}\geq 4$, $M_{3}\geq 6$ and $M_{4}\geq 1$ respectively.
@@
 which is a contradiction.
-Suppose that $\gamma=6$, then $b\leq\frac{7}{2}$. This contradicts to is contradictory to $b>\gamma-3$.
-We have completed the proof. ∎
-To conclude this section, we prove the number theoretic result Corollary 1.5.
+Suppose that $\gamma=6$, then $b\leq\frac{7}{2}$, which is contradictory to $b>\gamma-3$. ∎
+To conclude this section, we prove the number-theoretic Corollary 1.5.
 Proof of Corollary 1.5.
@@
 Since $|S|\geq\frac{3}{4}(n-1)$, we have $\left|\left(GF(n)\setminus\{x,y\}\right)\setminus S\right|\leq\frac{1}{4}(n-1)-1$.
-Recall a Paley graph $P(n)$ with $n=4\gamma+1$ is a strongly regular graph with parameters $(4\gamma+1,2\gamma,\gamma-1,\gamma)$.
+Recall that a Paley graph $P(n)$ with $n=4\gamma+1$ is a strongly regular graph with parameters $(4\gamma+1,2\gamma,\gamma-1,\gamma)$.
 Therefore, we have $|N_{x}|=\gamma=\frac{1}{4}(n-1)$.
 That is, we have $\left|\left(GF(n)\setminus\{x,y\}\right)\setminus S\right|\leq|N_{x}|-1$.
-Then it is clear that $GF(n)\backslash S$ is not a vertex cover of $\mathcal{M}$.
-Suppose that $wz\in\mathcal{M}$ is not covered by $GF(n)\setminus S$, then the result follows.
-∎
-4 Proof of Theorem 1.7 and Corollary 1.8
-In this section, we prove Theorem 1.7 and Corollary 1.8. We first recall the following facts about eigenvalues of strongly regular graphs, see, e.g., [17, Page 220].
-Theorem 4.1.
-Let $G=(V,E)$ be a strongly regular graph with parameters $(n,d,\alpha,\beta)$. Then the Laplacian eigenvalues of $G$ are
-$\lambda_{1}=0,$
-$\lambda_{2}=1-\frac{(\alpha-\beta)+\sqrt{(\alpha-\beta)^{2}+4(d-\beta)}}{2d},$
-$\lambda_{3}=1-\frac{(\alpha-\beta)-\sqrt{(\alpha-\beta)^{2}+4(d-\beta)}}{2d}.$
-The corresponding multiplicities are
-$m_{1}=1,$
-$m_{2}=\frac{1}{2}\biggl{(}(n-1)-\frac{2d+(n-1)(\alpha-\beta)}{\sqrt{(\alpha-\beta)^{2}+4(d-\beta)}}\biggr{)},$
-$m_{3}=\frac{1}{2}\biggl{(}(n-1)+\frac{2d+(n-1)(\alpha-\beta)}{\sqrt{(\alpha-\beta)^{2}+4(d-\beta)}}\biggr{)}.$
-By Theorem 1.1, Theorem 1.7 is a direct consequence of the following result.
-Theorem 4.2.
-Let $G=(V,E)$ be a strongly regular graph with parameters $(n,d,\alpha,\beta)$ such that $\beta\geq\alpha$.
-If $\lambda_{2}\leq\frac{2+\alpha}{d}$, then the equality holds, i.e., $\lambda_{2}=\frac{2+\alpha}{d}$.
-Proof.
-Since $\lambda_{2}\leq\frac{2+\alpha}{d}$, we have by Theorem 4.1 that
-$2d-3\alpha+\beta-4\leq\sqrt{(\alpha-\beta)^{2}+4(d-\beta)}.$ (4.1)
-Since $G$ is connected and not complete, we have $d\geq\alpha+2$. Therefore, we derive
-$2d-3\alpha+\beta-4\geq 2(\alpha+2)-3\alpha+\beta-4\geq 0$
-In the second inequality above, we use our assumption $\beta\geq\alpha$. It follows that we can square both sides of the inequality (4.1). Hence, we have
-$(2d-3\alpha+\beta-4)^{2}\leq(\alpha-\beta)^{2}+4(d-\beta).$
-The above inequality can be rearranged as
-$4(d-\alpha-1)(d-2\alpha+\beta-4)\leq 0$
-Since $d\geq\alpha+2$, we deduce
-$d\leq 2\alpha-\beta+4.$ (4.2)
-On one hand, since $\beta\geq\alpha$, we have $d\leq\alpha+4$. That is, $\alpha+2\leq d\leq\alpha+4$. On the other hand, since $d\geq\beta$, we have $\beta\leq\alpha+2$.
-Observe that the multiplicity numbers $m_{2}$ and $m_{3}$ in Theorem 4.1 must be positive integers.
-Hence, $(\alpha-\beta)^{2}+4(d-\beta)$ has to be a perfect square.
-Then, it is straightforward to check that there are only three possible parameters $(n,d,\alpha,\beta)$ for $G$:
-  • (a)
-    $(n,\alpha+4,\alpha,\alpha)$;
-  • (b)
-    $(n,\alpha+3,\alpha,\alpha+1)$;
-  • (c)
-    $(n,\alpha+2,\alpha,\alpha+2)$.
-Notice that $n$ can be uniquely determined by the other parameters $d,\alpha,\beta$. Indeed, we have by Theorem 2.6 that
-$n-d-1=\frac{d(d-\alpha-1)}{\beta}.$ (4.3)
-In the case of $(n,\alpha+4,\alpha,\alpha)$.
-Since $\frac{d(d-\alpha-1)}{\beta}=3+\frac{12}{\alpha}$ must be a positive integer, the possible values of $\alpha$ are $1$, $2$, $3$, $4$, $6$, $12$.
-By (4.3), we have $n=\alpha+\frac{12}{\alpha}+8$ and hence the multiplicity numbers $m_{2}=\frac{(4+\alpha)(6+\alpha)}{4\alpha}$ and $m_{3}=\frac{3(2+\alpha)(4+\alpha)}{4\alpha}$.
-Since $m_{2}$ and $m_{3}$ have to be positive integers, the possible values of $\alpha$ are $2$, $4$, $6$, $12$.
-Therefore, the possible parameters in this case are
-  • (1)
-    $(16,6,2,2)$;
-  • (2)
-    $(15,8,4,4)$;
-  • (3)
-    $(16,10,6,6)$;
-  • (4)
-    $(21,16,12,12)$.
-In the case of $(n,\alpha+3,\alpha,\alpha+1)$.
-Since $\frac{d(d-\alpha-1)}{\beta}=2+\frac{4}{\alpha+1}$ must be a positive integer, the possible values of $\alpha$ are 00, $1$, $3$.
-Therefore, the possible parameters are
-  • (5)
-    $(10,3,0,1)$;
-  • (6)
-    $(9,4,1,2)$;
-  • (7)
-    $(10,6,3,4)$.
-In the case of $(n,\alpha+2,\alpha,\alpha+2)$.
-By (4.3) and Theorem 4.1, we calculate that $m_{2}=\frac{4+\alpha}{2}$ and $m_{3}=\frac{2+\alpha}{2}$.
-Since $m_{2}$ and $m_{3}$ have to be positice integers, $\alpha$ must be even.
-Therefore, the possible parameters are
-  • (8)
-    $(2k,2k-2,2k-4,2k-2),k\geq 2$.
-By Theorem 4.1, it is direct to check for all possible graphs $(1)-(8)$ that
-$\lambda_{2}=\frac{2+\alpha}{d}.$
-This completes the proof. ∎
-Proof of Theorem 1.7.
-Since $G$ is Lichnerowicz sharp, we have
-$\lambda_{2}=\min_{xy\in E}\kappa(x,y)\leq\frac{2+\alpha}{d},$
-where the last inequality is due to Theorem 1.1.
-By Theorem 4.2, we conclude that $\min_{xy\in E}\kappa(x,y)=\lambda_{2}=\frac{2+\alpha}{d}$.
-Therefore, we obtain $\kappa(x,y)=\frac{2+\alpha}{d}$ for all $xy\in E$ by applying Theorem 1.1 again.
-This completes the proof.
-∎
-We recall the following classificaiton result from [14].
-Theorem 4.3.
-[14, Theorems 6.3 and 6.4] The Lichnerowicz sharp strongly regular graphs with smallest adjacency eigenvalue $-2$ are precisely the following ones: The cocktail party graphs $CP(n)$, $n\geq 2$, the lattice graphs $L_{2}(n)\cong K_{n}\times K_{n}$, $n\geq 3$, the triangular graphs $T(n)\cong J(n,2)$, $n\geq 5$, the demi-cube $Q^{5}_{(2)}$, and the Schläfli graph.
-Proof of Corollary 1.8.
-By Theorem 4.1, any strongly regular graph with $\lambda_{2}=\frac{2+\alpha}{d}$ has the smallest adjacency eigenvalue $-2$.
-Thus, by Theorem 4.3, we determine all the Lichnerowicz sharp strongly regular graphs with $\beta\geq\alpha$ as following: The cocktail party graphs $CP(k),k\geq 2$, the Paley graph $P(9)$, the Johnson graphs $J(5,2)$ and $J(6,2)$ and the demi-cube $Q^{5}_{(2)}$.
+Then, it is clear that there exists an edge $wz\in\mathcal{M}$, which is not adjacent to any vertex in $GF(n)\setminus S$, and the desired result follows.
 ∎
 Acknowledgement
-This work has been supported by the National Key R & D Program of China 2023YFA1010200 and the National Natural Science Foundation of China No.
+This work is supported by the National Key R & D Program of China 2023YFA1010200 and the National Natural Science Foundation of China No.
 12031017 and No.
@@
 is very grateful to Professor Wen Huang for his encouragement and helpful discussion.
-Appendix
-In this appendix, we present two lists of parameters $(n,d,\alpha,\beta)$ for strongly regular graphs with at most $512$ vertices based on the data in [9, Chapter 12]: Table LABEL:table:1234 provides strongly regular graphs in the list of [9, Chapter 12] that fulfill at least one of the conditions (1), (2), (3), and (4), and particularly, $\beta>\alpha$; Table LABEL:table:5 includes strongly regular graphs in the list of [9, Chapter 12] that meet the condition (5), and particularly $\beta\leq\alpha$.
-We first explain the notations column by column following [9, Chapter 12].
-  • (i)
-    Existence: A number indicates the precise number of non-isomorphic examples. "!" when there is a unique such graph, "+" when there is a known example, and "?" when there is not a known example.
-  • (ii)
-    The parameters $(n,d,\alpha,\beta)$: the number of vertices, the degree, the number of common neighbors of two adjacent vertices, and the number of common neighbors of two nonadjacent vertices, respectively.
-  • (iii)
-    The smallest non-zero Laplacian eigenvalue $\lambda_{2}$ and the Lin-Lu-Yau curvature $\kappa$.
-  • (iv)
-    Comments.
-The symbol ‘↓’ labels a descendant of a regular $2$-graph on $v+1$ vertices.
-The symbol ‘↑’ labels a graph in the switching class of a regular $2$-graph.
-Labels are postfixed with ‘?’ when the corresponding object is unknown.
-Table 1: Some strongly regular graphs satisfying one of conditions (1)-(4)
-ex n d $\alpha$ $\beta$ $\lambda_{2}$ $k(x,y)$ comment
-15! 25 12 5 6 $\frac{5}{6}$ $\frac{7}{12}$ Paulus-Rozenfel’d; OA(5, 3); Paley(25); ↓
-41! 29 14 6 7 $\frac{1181}{1400}$ $\frac{4}{7}$ Enumerated by Bussemaker and by Spence; Paley(29); ↓
-3854! 35 16 6 8 $\frac{7}{8}$ $\frac{1}{2}$ Enumerated by McKay & Spence; no pg(5, 4, 2) (De Clerck); ↓
-+ 37 18 8 9 $\frac{773}{900}$ $\frac{5}{9}$ see McKay & Spence; Crnković & Maksimović ; Maksimović & Rukavina; Paley(37); ↓
-+ 41 20 9 10 $\frac{173}{200}$ $\frac{11}{20}$ Maksimović & Rukavina; Paley(41);
-+ 45 22 10 11 $\frac{383}{440}$ $\frac{6}{11}$ Mathon; ↓
-+ 49 24 11 12 $\frac{7}{8}$ $\frac{13}{24}$ Paley(49); OA(7, 4); ↓
-+ 50 28 15 16 $\frac{25}{28}$ $\frac{17}{28}$ S(2, 4, 25); ↑
-+ 53 26 12 13 $\frac{1143}{1300}$ $\frac{7}{13}$ Paley(53); ↓
-+ 61 30 14 15 $\frac{2659}{3000}$ $\frac{8}{15}$ Paley(61); ↓
-+ 63 30 13 15 $\frac{9}{10}$ $\frac{1}{2}$ qs $2-(36,16,12)$; $O_{7}(2)$; $Sp_{6}(2)$; pg(7, 5, 3); ↓
-+ 64 35 18 20 $\frac{32}{35}$ $\frac{4}{7}$ OA(8, 5); Wallis; Goethals-Seidel; $VO^{+}_{6}(2)$; ↑
-+ 65 32 15 16 $\frac{2847}{3200}$ $\frac{17}{32}$ Gritsenko; ↓
-+ 73 36 17 18 $\frac{3223}{3600}$ $\frac{19}{36}$ Paley(73); ↓
-+ 81 40 19 20 $\frac{9}{10}$ $\frac{21}{40}$ Paley(81); OA(9, 5); ↓
-+ 82 45 24 25 $\frac{41}{45}$ $\frac{26}{45}$ S(2, 5, 41); ↑
-? 85 42 20 21 $\frac{1263}{1400}$ $\frac{11}{21}$
-+ 89 44 21 22 $\frac{1989}{2200}$ $\frac{23}{44}$ Paley(89); ↓
-+ 97 48 23 24 $\frac{2179}{2400}$ $\frac{25}{48}$ Paley(97); ↓
-+ 99 48 22 24 $\frac{11}{12}$ $\frac{1}{2}$ no pg(9, 6, 4)(Lam et al.); ↓
-+ 100 54 28 30 $\frac{25}{27}$ $\frac{5}{9}$ OA(10, 6)?; ↑
-+ 101 50 24 25 $\frac{1137}{1250}$ $\frac{13}{25}$ Paley(101); ↓
-? 105 64 38 40 $\frac{15}{16}$ $\frac{5}{8}$
-+ 109 54 26 27 $\frac{616}{675}$ $\frac{14}{27}$ Paley(109); ↓
-+ 113 56 27 28 $\frac{2559}{2800}$ $\frac{29}{56}$ Paley(113); ↓
-? 117 58 28 29 $\frac{5309}{5800}$ $\frac{15}{29}$?
-? 120 84 58 60 $\frac{20}{21}$ $\frac{5}{7}$
-+ 121 60 29 30 $\frac{11}{12}$ $\frac{31}{60}$ Paley(121); OA(11, 6); ↓
-+ 122 66 35 36 $\frac{61}{66}$ $\frac{37}{66}$ S(2, 6, 61); ↑
-+ 125 52 15 26 $\frac{25}{26}$ $\frac{17}{52}$ pg(5,13,2)?; ↓
-+ 125 62 30 31 $\frac{5691}{6200}$ $\frac{16}{31}$ Paley(125); ↓
-+ 137 68 33 34 $\frac{1253}{1360}$ $\frac{35}{68}$ Paley(137); ↓
-+ 143 70 33 35 $\frac{13}{14}$ $\frac{1}{2}$ qs $2-(78,36,30)$; pg(11,7,5)?; ↓
-+ 144 77 40 42 $\frac{72}{77}$ $\frac{6}{11}$ OA(12, 7); Wallis; Goethals-Seidel;
-? 145 72 35 36 $\frac{277}{300}$ $\frac{37}{72}$?
-+ 149 74 36 37 $\frac{171}{185}$ $\frac{19}{37}$ Paley(149); ↓
-? 153 76 37 38 $\frac{879}{950}$ $\frac{39}{76}$?
-? 154 72 26 40 $\frac{35}{36}$ $\frac{7}{18}$
-+ 157 78 38 39 $\frac{301}{325}$ $\frac{20}{39}$ Paley(157); ↓
-? 160 105 68 70 $\frac{20}{21}$ $\frac{2}{3}$
-+ 169 84 41 42 $\frac{13}{14}$ $\frac{43}{84}$ Paley(169); OA(13,7); ↓
-+ 170 91 48 49 $\frac{85}{91}$ $\frac{50}{91}$ S(2, 7, 85);
-+ 173 86 42 43 $\frac{999}{1075}$ $\frac{22}{43}$ Paley(173); ↓
-+ 175 72 20 36 $\frac{35}{36}$ $\frac{11}{36}$ Haemers pg(5,18,2); edges of Hoffman-Singleton; ↓
-+ 176 90 38 54 $\frac{44}{45}$ $\frac{4}{9}$ pg(6,18,3)?; ↑
-+ 181 90 44 45 $\frac{8377}{9000}$ $\frac{23}{45}$ Paley(181); ↓
-? 185 92 45 46 $\frac{857}{920}$ $\frac{47}{92}$?
-? 189 140 103 105 $\frac{27}{28}$ $\frac{3}{4}$
-+ 193 96 47 48 $\frac{597}{640}$ $\frac{49}{96}$ Paley(193); ↓
-+ 195 96 46 48 $\frac{15}{16}$ $\frac{1}{2}$ pg(13,8,6)?; ↓
-+ 196 114 59 76 $\frac{56}{57}$ $\frac{61}{114}$ pg(7,19,4)?
-+ 196 104 54 56 $\frac{49}{52}$ $\frac{7}{13}$ OA(14,8)?; ↑
-+ 197 98 48 49 $\frac{2287}{2450}$ $\frac{25}{49}$ Paley(197); ↓
-? 205 102 50 51 $\frac{1589}{1700}$ $\frac{26}{51}$?
-? 210 132 82 84 $\frac{21}{22}$ $\frac{7}{11}$
-? 221 110 54 55 $\frac{937}{1000}$ $\frac{28}{55}$?
-+ 225 112 55 56 $\frac{15}{16}$ $\frac{57}{112}$ ConfMat$(16)^{2}$; OA(15, 8)?; ↓
-+ 226 120 63 64 $\frac{113}{120}$ $\frac{13}{24}$ S(2, 8, 113);
-+ 229 114 56 57 $\frac{10693}{11400}$ $\frac{29}{57}$ Paley(229); ↓
-? 231 160 110 112 $\frac{77}{80}$ $\frac{7}{10}$
-+ 233 116 57 58 $\frac{10887}{11600}$ $\frac{59}{116}$ Paley(233); ↓
-+ 241 120 59 60 $\frac{1879}{2000}$ $\frac{61}{120}$ Paley(241); ↓
-+ 243 110 37 60 $\frac{54}{55}$ $\frac{39}{110}$ Delsarte; $[55,5]_{3}$ (wts 36,45)
-? 245 108 39 54 $\frac{35}{36}$ $\frac{41}{108}$ pg(7,18,3)?; ↓?
-? 245 122 60 61 $\frac{11467}{12200}$ $\frac{31}{61}$?
-+ 253 112 36 60 $\frac{55}{56}$ $\frac{19}{56}$ S(4, 7, 23); $M_{23}$
-+ 255 126 61 63 $\frac{17}{18}$ $\frac{1}{2}$ $O_{9}(2)$; $Sp_{8}(2)$; pg(15, 9, 7); ↓
-+ 256 135 70 72 $\frac{128}{135}$ $\frac{8}{15}$ OA(16, 9); Wallis; Goethals-Seidel; $VO^{+}_{8}(2)$; ↑
-+ 257 128 63 64 $\frac{753}{800}$ $\frac{65}{128}$ Paley(257); ↓
-? 261 130 64 65 $\frac{6121}{6500}$ $\frac{33}{65}$?
-? 265 132 65 66 $\frac{3109}{3300}$ $\frac{67}{132}$?
-+ 269 134 66 67 $\frac{1263}{1340}$ $\frac{34}{67}$ Paley(269); ↓
-! 275 112 30 56 $\frac{55}{56}$ $\frac{2}{7}$ $q_{22}^{2}=0$; no pg(5, 28, 2)
-+ 276 140 58 84 $\frac{69}{70}$ $\frac{3}{7}$ pg(6,28,3)?; ↑
-+ 277 138 68 69 $\frac{283}{300}$ $\frac{35}{69}$ Paley(277); ↓
-? 280 216 166 168 $\frac{35}{36}$ $\frac{7}{9}$
-+ 281 140 69 70 $\frac{3303}{3500}$ $\frac{71}{140}$ Paley(281); ↓
-? 287 126 45 63 $\frac{41}{42}$ $\frac{47}{126}$ pg(7,21,3)?; ↓?
-? 288 182 106 130 $\frac{90}{91}$ $\frac{54}{91}$ pg(8,26,5)?
-? 288 147 66 84 $\frac{48}{49}$ $\frac{68}{147}$ pg(8,21,4)?; ↑?
-+ 289 144 71 72 $\frac{17}{18}$ $\frac{73}{144}$ Paley(289); OA(17,9); ↓
-+ 290 153 80 81 $\frac{145}{153}$ $\frac{82}{153}$ S(2,9,145)?; ↑
-+ 293 146 72 73 $\frac{6897}{7300}$ $\frac{37}{73}$ Paley(293); ↓
-? 300 182 100 126 $\frac{90}{91}$ $\frac{51}{91}$
-? 301 150 65 84 $\frac{49}{50}$ $\frac{67}{150}$
-? 305 152 75 76 $\frac{14377}{15200}$ $\frac{77}{152}$?
-+ 313 156 77 78 $\frac{2953}{3120}$ $\frac{79}{156}$ Paley(313); ↓
-+ 317 158 78 79 $\frac{374}{395}$ $\frac{40}{79}$ Paley(317); ↓
-? 320 231 166 168 $\frac{32}{33}$ $\frac{8}{11}$
-+ 323 160 78 80 $\frac{19}{20}$ $\frac{1}{2}$ pg(17,10,8)?; ↓
-+ 324 170 88 90 $\frac{81}{85}$ $\frac{9}{17}$ OA(18,10)?; ↑
-? 325 252 195 196 $\frac{35}{36}$ $\frac{197}{252}$
-? 325 162 80 81 $\frac{15349}{16200}$ $\frac{41}{81}$?
-? 333 166 82 83 $\frac{7869}{8300}$ $\frac{42}{83}$?
-? 336 200 118 120 $\frac{24}{25}$ $\frac{3}{5}$
-+ 337 168 83 84 $\frac{569}{600}$ $\frac{85}{168}$ Paley(337); ↓
-+ 343 150 53 75 $\frac{49}{50}$ $\frac{11}{30}$ Godsil; pg(7,25,3)?; ↓
-+ 344 175 78 100 $\frac{172}{175}$ $\frac{16}{35}$ pg(8,25,4)?; Taylor ↑
-+ 349 174 86 87 $\frac{4129}{4350}$ $\frac{44}{87}$ Paley(349); ↓
-? 351 280 223 224 $\frac{39}{40}$ $\frac{45}{56}$
-+ 351 224 142 144 $\frac{27}{28}$ $\frac{9}{14}$ $NO_{7}^{-}(3)$
-? 351 210 113 144 $\frac{104}{105}$ $\frac{23}{42}$
-+ 353 176 87 88 $\frac{16711}{17600}$ $\frac{89}{176}$ Paley(353); ↓
-+ 361 180 89 90 $\frac{19}{20}$ $\frac{91}{180}$ Paley(361); OA(19,10); ↓
-+ 362 190 99 100 $\frac{181}{190}$ $\frac{101}{190}$ S(2,10,181)?; ↑
-? 365 182 90 91 $\frac{3459}{3640}$ $\frac{46}{91}$?
-? 369 184 91 92 $\frac{1749}{1840}$ $\frac{93}{184}$?
-+ 373 186 92 93 $\frac{4421}{4650}$ $\frac{47}{93}$ Paley(373); ↓
-? 375 198 93 117 $\frac{65}{66}$ $\frac{95}{198}$
-? 377 188 93 94 $\frac{17879}{18800}$ $\frac{95}{188}$?
-? 378 260 178 180 $\frac{63}{65}$ $\frac{9}{13}$ $NO_{7}^{+}(3)$
-+ 389 194 96 97 $\frac{2308}{2425}$ $\frac{49}{97}$ Paley(389); ↓
-+ 397 198 98 99 $\frac{857}{900}$ $\frac{50}{99}$ Paley(397); ↓
-+ 399 198 97 99 $\frac{21}{22}$ $\frac{1}{2}$ pg(19,11,9)?; ↓
-+ 400 209 108 110 $\frac{200}{209}$ $\frac{10}{19}$ OA(20,11)?; Wallis; Goethals-Seidel; ↑
-+ 401 200 99 100 $\frac{19049}{20000}$ $\frac{101}{200}$ Paley(401); ↓
-? 405 202 100 101 $\frac{4811}{5050}$ $\frac{51}{101}$?
-+ 409 204 101 102 $\frac{19439}{20400}$ $\frac{103}{204}$ Paley(409); ↓
-+ 421 210 104 105 $\frac{2503}{2625}$ $\frac{53}{105}$ Paley(421); ↓
-? 425 264 163 165 $\frac{85}{88}$ $\frac{5}{8}$
-? 425 212 105 106 $\frac{20219}{21200}$ $\frac{107}{212}$?
-? 429 320 238 240 $\frac{39}{40}$ $\frac{3}{4}$
-+ 433 216 107 108 $\frac{229}{240}$ $\frac{109}{216}$ Paley(433); ↓
-? 441 220 95 124 $\frac{217}{220}$ $\frac{97}{220}$
-+ 441 220 109 110 $\frac{21}{22}$ $\frac{111}{220}$ Mathon; OA(21,11)?; ↓
-+ 442 231 120 121 $\frac{221}{231}$ $\frac{122}{231}$ S(2,11,221)?; ↑?
-? 445 222 110 111 $\frac{1413}{1480}$ $\frac{56}{111}$?
-? 448 297 196 198 $\frac{32}{33}$ $\frac{2}{3}$
-+ 449 224 111 112 $\frac{21391}{22400}$ $\frac{113}{224}$ Paley(449); ↓
-+ 457 228 113 114 $\frac{21781}{22800}$ $\frac{115}{228}$ Paley(457); ↓
-? 459 208 82 104 $\frac{51}{52}$ $\frac{21}{52}$ pg(9,26,4)?; ↓?
-? 460 243 114 144 $\frac{80}{81}$ $\frac{116}{243}$
-+ 461 230 114 115 $\frac{2747}{2875}$ $\frac{58}{115}$ Paley(461); ↓
-? 477 238 118 119 $\frac{11379}{11900}$ $\frac{60}{119}$?
-? 481 240 119 120 $\frac{7651}{8000}$ $\frac{121}{240}$?
-? 483 240 118 120 $\frac{23}{24}$ $\frac{1}{2}$ pg(21,12,10)?; ↓?
-? 484 252 130 132 $\frac{121}{126}$ $\frac{11}{21}$ OA(22,12)?; ↑?
-? 485 242 120 121 $\frac{23149}{24200}$ $\frac{61}{121}$?
-? 493 246 122 123 $\frac{1177}{1230}$ $\frac{62}{123}$?
-? 495 234 93 126 $\frac{77}{78}$ $\frac{95}{234}$
-? 505 252 125 126 $\frac{4021}{4200}$ $\frac{127}{252}$?
-? 507 276 135 168 $\frac{91}{92}$ $\frac{137}{276}$
-+ 509 254 126 127 $\frac{12161}{12700}$ $\frac{64}{127}$ Paley(509); ↓
-Table 2: Some strongly regular graphs satisfying the condition (5)
-ex n d $\alpha$ $\beta$ $\lambda_{2}$ $k(x,y)$ comment
-28! 40 27 18 18 $\frac{8}{9}$ $\frac{20}{27}$ $NU_{4}(2)$
-! 56 45 36 36 $\frac{14}{15}$ $\frac{38}{45}$ qs 2-(21,6,4)
-167! 64 45 32 30 $\frac{8}{9}$ $\frac{34}{45}$
-! 77 60 47 45 $\frac{11}{12}$ $\frac{49}{60}$ Witt: qs 2-(22,6,5)
-! 81 60 45 42 $\frac{9}{10}$ $\frac{47}{60}$
-+ 85 64 48 48 $\frac{15}{16}$ $\frac{25}{32}$
-? 88 60 41 40 $\frac{11}{12}$ $\frac{43}{60}$
-+ 96 76 60 60 $\frac{18}{19}$ $\frac{31}{38}$
-! 100 77 60 56 $\frac{10}{11}$ $\frac{62}{77}$ $q_{11}^{1}=0$
-? 111 80 58 56 $\frac{37}{40}$ $\frac{3}{4}$
-! 112 81 60 54 $\frac{8}{9}$ $\frac{62}{81}$ $q_{11}^{1}=0$
-? 112 75 50 50 $\frac{14}{15}$ $\frac{52}{75}$
-? 115 96 80 80 $\frac{23}{24}$ $\frac{41}{48}$
-? 120 85 60 60 $\frac{16}{17}$ $\frac{62}{85}$
-? 121 84 59 56 $\frac{11}{12}$ $\frac{61}{84}$
-+ 125 96 74 72 $\frac{15}{16}$ $\frac{19}{24}$
-? 133 100 75 75 $\frac{19}{20}$ $\frac{77}{100}$
-+ 144 104 76 72 $\frac{12}{13}$ $\frac{3}{4}$
-? 154 105 72 70 $\frac{14}{15}$ $\frac{74}{105}$
-+ 156 125 100 100 $\frac{24}{25}$ $\frac{102}{125}$
-? 162 140 121 120 $\frac{27}{28}$ $\frac{123}{140}$
-+ 162 128 100 96 $\frac{15}{16}$ $\frac{51}{64}$
-? 169 126 95 90 $\frac{13}{14}$ $\frac{97}{126}$
-? 171 120 84 84 $\frac{19}{20}$ $\frac{43}{60}$
-? 176 150 128 126 $\frac{24}{25}$ $\frac{13}{15}$
-+ 176 126 90 90 $\frac{20}{21}$ $\frac{46}{63}$
-? 183 130 93 90 $\frac{61}{65}$ $\frac{19}{26}$
-? 196 156 125 120 $\frac{49}{52}$ $\frac{127}{156}$
-? 196 150 116 110 $\frac{14}{15}$ $\frac{59}{75}$
-+ 196 135 94 90 $\frac{14}{15}$ $\frac{32}{45}$
-? 204 175 150 150 $\frac{34}{35}$ $\frac{152}{175}$
-? 208 162 126 126 $\frac{26}{27}$ $\frac{64}{81}$
-? 210 176 148 144 $\frac{21}{22}$ $\frac{75}{88}$
-+ 216 175 142 140 $\frac{24}{25}$ $\frac{144}{175}$
-? 217 150 105 100 $\frac{14}{15}$ $\frac{107}{150}$
-? 220 147 98 98 $\frac{20}{21}$ $\frac{100}{147}$
-? 225 176 139 132 $\frac{15}{16}$ $\frac{141}{176}$
-? 225 160 115 110 $\frac{15}{16}$ $\frac{117}{160}$
-? 232 198 169 168 $\frac{32}{33}$ $\frac{19}{22}$
-? 232 168 122 120 $\frac{20}{21}$ $\frac{31}{42}$
-? 235 182 141 140 $\frac{25}{26}$ $\frac{11}{14}$
-? 238 162 111 108 $\frac{17}{18}$ $\frac{113}{162}$
-? 243 176 130 120 $\frac{81}{88}$ $\frac{3}{4}$
-? 245 192 152 144 $\frac{15}{16}$ $\frac{77}{96}$
-? 250 168 113 112 $\frac{20}{21}$ $\frac{115}{168}$
-+ 256 204 164 156 $\frac{16}{17}$ $\frac{83}{102}$ vanLint-Schrijver
-+ 256 187 138 132 $\frac{16}{17}$ $\frac{140}{187}$
-? 259 216 180 180 $\frac{35}{36}$ $\frac{91}{108}$
-? 260 189 138 135 $\frac{20}{21}$ $\frac{20}{27}$
-? 261 196 147 147 $\frac{27}{28}$ $\frac{149}{196}$
-? 266 220 183 176 $\frac{19}{20}$ $\frac{37}{44}$
-? 273 192 136 132 $\frac{91}{96}$ $\frac{23}{32}$
-? 276 200 148 136 $\frac{23}{25}$ $\frac{3}{4}$
-? 276 200 145 144 $\frac{24}{25}$ $\frac{147}{200}$
-? 280 217 168 168 $\frac{30}{31}$ $\frac{170}{217}$
-? 285 220 171 165 $\frac{19}{20}$ $\frac{173}{220}$
-? 288 246 210 210 $\frac{40}{41}$ $\frac{106}{123}$
-? 289 216 163 156 $\frac{17}{18}$ $\frac{55}{72}$
-? 289 198 137 132 $\frac{17}{18}$ $\frac{139}{198}$
-+ 300 234 183 180 $\frac{25}{26}$ $\frac{185}{234}$
-? 306 250 205 200 $\frac{24}{25}$ $\frac{207}{250}$
-? 306 245 196 196 $\frac{34}{35}$ $\frac{198}{245}$
-? 320 232 168 168 $\frac{28}{29}$ $\frac{85}{116}$
-? 322 225 160 150 $\frac{14}{15}$ $\frac{18}{25}$
-? 324 255 202 195 $\frac{81}{85}$ $\frac{4}{5}$
-? 324 247 190 182 $\frac{18}{19}$ $\frac{192}{247}$
-? 324 228 162 156 $\frac{18}{19}$ $\frac{41}{57}$
-? 325 270 225 220 $\frac{26}{27}$ $\frac{227}{270}$
-+ 325 256 204 192 $\frac{15}{16}$ $\frac{103}{128}$ $q_{11}^{1}=0$
-? 329 288 252 252 $\frac{47}{48}$ $\frac{127}{144}$
-? 340 231 158 154 $\frac{20}{21}$ $\frac{160}{231}$
-? 341 256 192 192 $\frac{31}{32}$ $\frac{97}{128}$
-+ 343 288 242 240 $\frac{35}{36}$ $\frac{61}{72}$
-? 343 240 171 160 $\frac{14}{15}$ $\frac{173}{240}$
-? 352 325 300 300 $\frac{64}{65}$ $\frac{302}{325}$
-? 352 315 282 280 $\frac{44}{45}$ $\frac{284}{315}$
-? 361 280 219 210 $\frac{19}{20}$ $\frac{221}{280}$
-? 361 260 189 182 $\frac{19}{20}$ $\frac{191}{260}$
-? 364 275 210 200 $\frac{52}{55}$ $\frac{212}{275}$
-+ 364 243 162 162 $\frac{26}{27}$ $\frac{164}{243}$
-? 375 308 253 252 $\frac{75}{77}$ $\frac{255}{308}$
-? 375 264 188 180 $\frac{125}{132}$ $\frac{95}{132}$
-+ 378 325 280 275 $\frac{63}{65}$ $\frac{282}{325}$
-+ 378 261 180 180 $\frac{28}{29}$ $\frac{182}{261}$
-? 381 266 187 182 $\frac{127}{133}$ $\frac{27}{38}$
-? 385 324 273 270 $\frac{35}{36}$ $\frac{275}{324}$
-? 392 345 304 300 $\frac{112}{115}$ $\frac{102}{115}$
-+ 400 343 294 294 $\frac{48}{49}$ $\frac{296}{343}$
-? 400 315 250 240 $\frac{20}{21}$ $\frac{4}{5}$
-? 400 294 218 210 $\frac{20}{21}$ $\frac{110}{147}$
-? 400 273 188 182 $\frac{20}{21}$ $\frac{190}{273}$
-? 405 308 235 231 $\frac{27}{28}$ $\frac{237}{308}$
-? 408 297 216 216 $\frac{32}{33}$ $\frac{218}{297}$
-? 428 315 234 225 $\frac{20}{21}$ $\frac{236}{315}$
-? 430 294 203 196 $\frac{20}{21}$ $\frac{205}{294}$
-? 437 336 260 252 $\frac{23}{24}$ $\frac{131}{168}$
-? 441 352 283 272 $\frac{21}{22}$ $\frac{285}{352}$
-? 441 330 249 240 $\frac{21}{22}$ $\frac{251}{330}$
-? 441 308 217 210 $\frac{21}{22}$ $\frac{219}{308}$
-? 451 320 228 224 $\frac{77}{80}$ $\frac{23}{32}$
-? 456 375 310 300 $\frac{24}{25}$ $\frac{104}{125}$
-? 456 351 270 270 $\frac{38}{39}$ $\frac{272}{351}$
-? 456 325 236 220 $\frac{304}{325}$ $\frac{238}{325}$
-? 456 315 218 216 $\frac{304}{315}$ $\frac{44}{63}$
-? 460 360 282 280 $\frac{35}{36}$ $\frac{71}{90}$
-? 460 312 213 208 $\frac{23}{24}$ $\frac{215}{312}$
-? 465 320 220 220 $\frac{31}{32}$ $\frac{111}{160}$
-? 470 343 252 245 $\frac{47}{49}$ $\frac{254}{343}$
-? 477 336 240 228 $\frac{53}{56}$ $\frac{121}{168}$
-? 484 391 318 306 $\frac{22}{23}$ $\frac{320}{391}$
-? 484 378 297 288 $\frac{121}{126}$ $\frac{299}{378}$
-? 484 368 282 272 $\frac{22}{23}$ $\frac{71}{92}$
-? 484 345 248 240 $\frac{22}{23}$ $\frac{50}{69}$
-? 486 388 310 308 $\frac{189}{194}$ $\frac{78}{97}$
-? 490 345 248 230 $\frac{14}{15}$ $\frac{50}{69}$
-? 494 408 337 336 $\frac{133}{136}$ $\frac{113}{136}$
-? 495 456 420 420 $\frac{75}{76}$ $\frac{211}{228}$
-? 496 441 392 392 $\frac{62}{63}$ $\frac{394}{441}$
-? 496 385 300 294 $\frac{372}{385}$ $\frac{302}{385}$
-? 505 420 351 340 $\frac{101}{105}$ $\frac{353}{420}$
-? 506 405 324 324 $\frac{44}{45}$ $\frac{326}{405}$
-? 507 462 421 420 $\frac{65}{66}$ $\frac{141}{154}$
-? 507 352 246 240 $\frac{169}{176}$ $\frac{31}{44}$
-? 511 432 366 360 $\frac{35}{36}$ $\frac{23}{27}$
-+ 512 441 380 378 $\frac{48}{49}$ $\frac{382}{441}$
-+ 512 378 282 270 $\frac{20}{21}$ $\frac{142}{189}$